### v1

# Weyl distributions, spectral properties, and circulant approximation results for quaternion block multilevel Toeplitz matrix sequences

Abstract
The present work contains a comprehensive treatment of Weyl eigenvalue and singular value distributions for single-axis quaternion block multilevel Toeplitz matrix sequences generated by $s\times t$ quaternion matrix-valued, $d$-variate, Lebesgue integrable generating functions. Furthermore, in view of concrete applications, we are interested in preconditioning and matrix approximation results. To this end, a crucial step is the extension of the notion of an approximating class of sequences (a.c.s.) to the case of matrix sequences with quaternion entries, since it allows us to decompose the difference between a matrix and its preconditioner into low-norm plus (relatively) low-rank terms. As a specific example, we consider classes of quaternion block multilevel circulant matrix sequences as an a.c.s. for quaternion block multilevel Toeplitz matrix sequences. These approximation results lay the foundations for fast preconditioning methods when dealing with large quaternion linear systems stemming from modern applications. We conclude our study with numerical experiments and directions for future research.
Keywords: Quaternions; Toeplitz matrices; matrix sequences; asymptotic spectral and singular value distributions; a.c.s. notion.
1 Introduction
The linear algebra and matrix theory of the quaternions [64] and their numerical applications have grown significantly in recent years; see, e.g., [25] for the foundations in the context of noncommutative algebras. In particular, structured matrices such as Toeplitz-type quaternion matrices are widely considered in applications such as discrete convolutions on tori and quaternion signal analysis (cf., for example, the motivations of the analysis in [35]), while structured quaternion matrix theory is emerging in fields such as color imaging, polarization imaging, and low-rank quaternion modeling; see, e.g., [27, 10, 42, 37] and the references therein. In these applications, there is a need for fast algorithms, as they are well-studied for the case of large linear systems with complex coefficient matrices. In particular, when Toeplitz structures are encountered, a successful strategy has consisted of approximating them using matrix algebras associated with fast transforms such as circulants, $\omega$-circulants, and trigonometric matrix algebras (see [39, 12, 47, 21, 5] and the references therein). In turn, these approximations are used in connection with Krylov methods, with the goal of designing effective preconditioners, or in connection with multigrid (multi-iterative) solvers for finding appropriate smoothers and projection operators. In this direction, it is worth emphasizing that there has been recent interest in designing structure-preserving quaternion Krylov methods (such as QGMRES, QBiCG, QFOM), for which both the modeling benefits and the algorithmic constraints are analyzed when working directly over $\mathbb{H}$ [26, 32, 33, 34].
On the other hand, the analysis of the efficiency of such preconditioners depends on an accurate spectral and singular value analysis of the original matrices and matrix sequences and of the preconditioned ones. In this direction, having in mind the Toeplitz setting in the complex case, we need precise localization results, a study of the extreme spectral/singular value behavior, and an asymptotic study of the Weyl eigenvalue and singular value distributions of Toeplitz matrix sequences with quaternion entries, using the specific features of quaternion matrix theory; see [64, 43]. The asymptotic and analytic study of the quaternion Toeplitz spectrum parallels the theory of canonical-eigenvalue computations [38] and modern quaternion low-rank/SVD analysis [57, 24].
More concretely, our aim is to develop an asymptotic spectral and singular value distribution theory for multilevel block quaternion Toeplitz sequences generated by Fourier transforms along a single-axis, and an approximation framework (that is, the approximating class of sequences (a.c.s.) theory) that helps to make the approximation theory of quaternion matrix sequences as systematic as in the complex case. Even though this step is far from trivial, we look at it as the first task for treating more general problems where there are locally varying coefficients. In fact, the goal is to build the quaternion counterpart of the systematic asymptotic spectral and singular value theory of the complex generalized locally Toeplitz (GLT) framework [20, 21, 5]. We believe that developing such theoretical machinery would represent a strong tool that could accelerate this emerging field.
The two main problems we face when developing quaternion spectral theory are well known: the noncommutativity of the quaternion algebra, which makes left and right actions genuinely different, and a theory of eigenvalues that is substantially different from the case of complex-valued matrices. Some intuitive solutions to these problems, which we adopt, are to analyze complex slices, $\mathbb{C}_{\mathbf{i}}$ separately and to use the “symplectic embedding” trick introduced by Lee [31] for many problems related to the properties of quaternions; see also [25]. Beyond the applications we suggest, quaternion models also appear in control, collective dynamics, discrete geometry, and computational number theory. This further motivates the development of a robust spectral and matrix theory over $\mathbb{H}$ [18, 11, 56, 28, 29].
In the current paper, we contribute in the following way:
  1. 1.
    We give a practical definition of eigenvalue and singular value symbols for quaternion sequences by analyzing the canonical eigenvalue, which reproduces most complex spectral asymptotic properties as given in [62, 59, 61, 13]. Canonical eigenvalues provide lossless representatives of all infinite right-eigenvalue classes; thus, we regard this choice as ideal for our purposes. The singular values of a quaternion matrix behave, in an asymptotic sense, in the same way as their complex matrix counterparts, so the setting we define is analogous. Since we consider block multilevel Toeplitz matrix sequences with $s\times t$ quaternion matrix-valued generating functions, the corresponding complex counterparts are represented by the works [59, 13]. We refer the reader to [50, 60] for the original definition and its first use, and to [20, Chapter 5] for the a.c.s. calculus in the complex setting.
  2. 2.
    We extend approximating classes of sequences to rectangular quaternionic matrix sequences and prove the equivalence of the defined approximation with complex a.c.s. after embedding, with matching rank and norm normalizations. As corollaries, we recover all the existing complex standard a.c.s. calculus over $\mathbb{H}$ (ACSH1–5), which is fundamental for the development of the asymptotic linear algebra of structured matrix sequences.
  3. 3.
    We formalize the three single-axis kernels and recover standard transport identities under the embedding for the rectangular matrix-valued case and the reduction to right kernels. In the scalar case, these are classical tools in the field of quaternion signal processing.
  4. 4.
    For multilevel block Toeplitz matrices we show that any sandwich Toeplitz matrix equals a right Toeplitz matrix with a reflected symbol; as a consequence, the corresponding matrix sequence shares the same property. Further, we explicitly prove adjoint identities and a sharp criterion for proving the Hermitian character in the block, multilevel setting, together with tight localization results for the associated spectra.
  5. 5.
    We give exact complex $2s\times 2t$ singular and canonical eigenvalue symbols for the left, right, and sandwich cases. We prove that multilevel quaternion Toeplitz sequences with $L^{1}$ symbols are distributed in singular values as the embedded symbols $G_{\tau}$ for $\tau\in\{L,R,(S_{L},S_{R})\}$. For eigenvalues, we treat (i) Hermitian sequences without additional regularity and (ii) non-Hermitian symbols under a Tilli-type separation hypothesis of the embedded generating function; for the separation hypothesis see [61, 13].
  6. 6.
    We recover the Schatten $p$-norm estimates for Toeplitz matrices with $L^{p}$ symbols in the same scaling as the complex case, with the exception of a factor $2$; see Theorem 4.13. For Hermitian sequences we localize spectra of all finite matrices inside the essential range of the embedded Hermitian symbol and we obtain positivity criteria. Here the reference in the complex setting is the paper [54], where a rich variety of variational and norm inequalities are derived.
  7. 7.
    Adapting the Pan-Ng splitting and decomposition [41] for our purposes and settings, we show that right Toeplitz sequences admit a circulant a.c.s., providing an alternative route to singular/spectral distribution results for quaternion Toeplitz sequences that recover the classical approach for the corresponding complex Toeplitz sequences.
In summary, our theoretical tools and matrix analysis results are based on [64, 43], with the canonical-eigenvalue selection from [64]. The symplectic embedding trick, to our knowledge, dates back to [31] and is here used at the level of matrix sequences. The use of a quaternion Fourier analysis in signal processing is discussed in [17, 16]. In the complex case, the multilevel GLT machinery is well developed and it is explained in an exhaustive way in the surveys [21, 5]. For the unilevel scalar Hermitian quaternion Toeplitz with bounded generating function, see [35].
Organization of the work.
In Section 2 we review algebraic primitives, symplectic embedding, canonical eigenvalues, Schatten norms, and function-space preliminaries. Section 3 contains the embedding-based notion of spectral and singular value symbols. In Section 4 we develop the study of single-axis quaternion Toeplitz sequences, we prove kernel reduction, adjoint and properties of quaternion Hermitian character, together with proper localization results of the spectra. Furthermore, we give exact embedded symbols and we derive distribution theorems and Schatten/localization bounds. In Section 4.6 we establish the circulant a.c.s. and the QDFT fiber picture, giving an alternative proof of the spectral and singular value Toeplitz distributions. Section 5 is devoted to numerical tests for verifying and visualizing distribution results and for checking whether they are already evident in practical and low dimensional contexts. Conclusions and further remarks are reported in Section 6, together with a coherent plan for future investigations.
2 Definitions and preliminary tools
In the present section we set the basic notation for quaternions and quaternion matrices, recall the slice algebras $\mathbb{C}_{\mu}$ and the spectral facts used later. For most of the notations, we follow the standard conventions for quaternion algebra and linear algebra. For a systematic and more theoretical treatment of quaternion linear algebra, see [43]. The link between quaternion and complex linear algebra is mostly taken from [64] and will be used repeatedly.
2.1 Algebraic primitives
The classical Hamilton quaternion algebra is the skew-field
$\mathbb{H}:=\operatorname{span}_{\mathbb{R}}\{1,\mathbf{i},\mathbf{j},\mathbf{k}\},\qquad\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=\mathbf{i}\mathbf{j}\mathbf{k}=-1,$
where, by construction,
$\mathbf{i}\mathbf{j}=\mathbf{k},\qquad\mathbf{j}\mathbf{k}=\mathbf{i},\qquad\mathbf{k}\mathbf{i}=\mathbf{j}.$
For simplicity in our calculations, in this paper, the unit $\mathbf{k}$ will always be represented in the explicit form $\mathbf{k}=\mathbf{i}\mathbf{j}$.
By construction $q\in\mathbb{H}$ has the Cartesian form $q=q_{0}+q_{1}\mathbf{i}+q_{2}\mathbf{j}+q_{3}\mathbf{k}=q_{0}+q_{1}\mathbf{i}+q_{2}\mathbf{j}+q_{3}\mathbf{i}\mathbf{j}$ with $q_{j}\in\mathbb{R}$ the real components. The component $\Re q:=q_{0}$ is called the real part, while $\Im q:=q_{1}\mathbf{i}+q_{2}\mathbf{j}+q_{3}\mathbf{k}$ is the imaginary part.
The quaternion conjugation and the modulus (or norm) are defined in a similar way as for complex numbers, that is, we have
$\overline{q}=q_{0}-q_{1}\mathbf{i}-q_{2}\mathbf{j}-q_{3}\mathbf{k},\qquad|q|:=\sqrt{q\,\overline{q}},$
so that |pq|=|p||q||pq|=|p|\,|q|, exactly as in the complex setting.
We call $\mu\in\mathbb{H}$ an imaginary unit if $\Re\mu=0$ and $\mu^{2}=-1$. Given such $\mu$, we can define the complex slice
$\mathbb{C}_{\mu}:=\operatorname{span}_{\mathbb{R}}\{1,\mu\}\cong\mathbb{C}.$
If $\nu$ is another normalized imaginary unit orthogonal to $\mu$, that means $\nu\mu=-\mu\nu$, then we find the important orthogonal decomposition (see [64, Thm. 2.1])
$\mathbb{H}\;=\;\mathbb{C}_{\mu}\oplus\mathbb{C}_{\mu}\nu,\qquad\text{i.e.}\quad q=z+w\,\nu\ \text{with}\ z,w\in\mathbb{C}_{\mu},$ (2.1)
and the representation in (2.1) is unique. Unless stated otherwise, when we perform an orthogonal slice decomposition, we fix for simplicity $\mu=\mathbf{i}$ and $\nu=\mathbf{j}$. The representation (2.1) is convenient for our matrix analysis, since it mimics complex algebra representations and enables a smooth use of complex-analytic arguments via the symplectic embedding defined in the next section.
In many algebraic arguments of this paper, we use the following elementary identities
$\mathbf{j}\alpha=\overline{\alpha}\,\mathbf{j}\quad\text{for all}\quad\alpha\in\mathbb{C}_{\mathbf{i}},$ (2.2)
and the associated matrix version
$\mathbf{j}A=\overline{A}\,\mathbf{j}\quad\text{for all}\quad A\in\mathbb{C}_{\mathbf{i}}^{m\times n}.$ (2.3)
For readability purposes, we declare explicitly these identities when we use them.
2.2 Symplectic embedding
We start by recalling the classical complex symplectic embedding. Writing $x=z+w\mathbf{j}$ with $z,w\in\mathbb{C}_{\mathbf{i}}$ as in (2.1) and taking into account the results in [31], define
$\Phi(x):=\begin{bmatrix}z&w\\ -\overline{w}&\overline{z}\end{bmatrix}\ \in\ \mathbb{C}^{2\times 2}.$ (2.4)
The definition of $\Phi$ can be extended to matrices as follows. If $A=Z+W\mathbf{j}$ with $Z,W\in\mathbb{C}_{\mathbf{i}}^{\,m\times n}$, then we set
$\Phi_{1}(A)\;:=\;\begin{bmatrix}Z&W\\ -\overline{W}&\overline{Z}\end{bmatrix}\ \in\ \mathbb{C}^{2m\times 2n}.$
Equivalently, one may apply (2.4) entrywise to every entry of a matrix. With a slight abuse of notation, we also denote the matrix entrywise extension by $\Phi$. The two extensions are related by a fixed permutation:
$P_{m}^{*}\,\Phi(A)\,P_{n}\;=\;\Phi_{1}(A),$ (2.5)
where $P_{n}\in\{0,1\}^{2n\times 2n}$ is the permutation
$P_{n}\;:=\;\sum_{i=1}^{n}\bigl(e_{2i-1}e_{i}^{*}+e_{2i}e_{n+i}^{*}\bigr),$
and $\{e_{j}\}_{j=1}^{2n}$ is the Euclidean basis of $\mathbb{C}^{2n}$.
The key structural property is indeed that $\Phi$ is a real-linear *-algebra embedding into complex matrices, see [31, 64]. Throughout, if either $X\in\mathbb{H}^{m\times t}$ or $\in\mathbb{C}^{m\times t}, XX^{*}$ denotes its transpose-conjugate, with respect to quaternions or complex numbers, respectively.
Lemma 2.1 (*-algebra property).
$\Phi$ is an injective real-linear *-algebra map. In particular,
$\Phi(AB)=\Phi(A)\Phi(B),\qquad\Phi(A^{*})=\Phi(A)^{*}.$
2.3 Spectral tools for quaternion matrices
Because of non-commutativity, left and right spectra differ in general and they have different properties. In this paper we always work with right eigenvalues, which together with their related eigenvectors, are defined as follows.
Definition 2.2 (Right eigenpair).
A nonzero $x\in\mathbb{H}^{m}$ and $\lambda\in\mathbb{H}$ form a right-eigenpair of $A\in\mathbb{H}^{m\times m}$ if $Ax=x\lambda$.
Right eigenvalues in general are infinitely many, but we can group them with quaternion conjugacy classes as reported in the next result. As a consequence, using these equivalence classes, we can manage the spectra of the considered matrix sequences as in the complex setting.
Lemma 2.3 (Conjugacy classes).
If $Ax=x\lambda$, then $A(xq)=(xq)(q^{-1}\lambda q)$ for every nonzero $q$. Thus right eigenvalues occur in conjugacy classes
$\lambda]:=\{\,q^{-1}\lambda q:q\neq 0\,\}.$
An important property is that, while eigenvalues are infinitely many, right conjugacy classes are not.
Lemma 2.4 ([64, Thm. 5.4]).
Every $A\in\mathbb{H}^{n\times n}$ has exactly $n$ right-eigenvalue classes, counted with multiplicity.
Note that, by Lemma 2.3, if an eigenvalue is real, then its conjugacy class collapses to the eigenvalue singleton. This is an important property of real eigenvalues that simplifies the analysis of Hermitian or normal matrices, for example.
Canonical eigenvalues.
We now recall a standard way to select meaningful representatives of right eigenvalue classes. This enables us to perform a complete matrix spectral analysis using a similar approach, as done in the complex setting. Indeed, Zhang proved the existence of the following complex canonical eigenvalues:
Theorem 2.5 ([64, Thm. 5.4]).
Every $A\in\mathbb{H}^{n\times n}$ has exactly $n$ right eigenvalues which are complex numbers with nonnegative imaginary parts, counted with multiplicity.
A simple way to compute canonical eigenvalues is to compute the spectrum of $\Phi(A)\in\mathbb{C}^{2n\times 2n}$. By construction, the eigenvalues of $\Phi(A)$ occur in conjugate pairs $\{z,\overline{z}\}$; if we select those with $\Im z\geq 0$ we obtain the $n$ (up to multiplicity) canonical eigenvalues. For subsequent applications, note that the selection of eigenvalues to $\{\Im z\geq 0\}$ is Borel measurable for conjugate pairs.
Quaternionic SVD and Schatten norms
The singular value decomposition (SVD) over $\mathbb{H}$ mirrors the complex case.
Theorem 2.6 (Quaternion SVD, [64, Thm. 7.2]).
For $A\in\mathbb{H}^{m\times n}$ there exist quaternionic unitary matrices $U,V$ of the proper dimensions, $r$ a nonnegative integer, and $\Sigma=\operatorname{diag}(\sigma_{1},\dots,\sigma_{r})$ with $\sigma_{j}>0$, $j=1,\ldots,r$, such that
$A=U\begin{bmatrix}\Sigma&0\\[1.0pt] 0&0\end{bmatrix}V^{*}.$
The $\sigma_{j}$ are the positive square roots of the positive eigenvalues of $A^{*}A$, the other eigenvalues being all equal to zero.
Definition 2.7 (Schatten $p$-norms).
Let $A\in\mathbb{H}^{m\times n}$ and set $r:=\min\{m,n\}$. Let the singular values be $\sigma_{1}(A)\geq\cdots\geq\sigma_{r}(A)\geq 0$. For $p\in[1,\infty)$ the Schatten $p$-norm of $A$ is
$\|A\|_{p}\;:=\;\Biggl(\,\sum_{j=1}^{r}\sigma_{j}(A)^{\,p}\,\Biggr)^{\!1/p},$
and for $p=\infty$ set $\|A\|_{\infty}:=\sigma_{1}(A)$. These norms are unitarily invariant over $\mathbb{H}$: $\|UAV\|_{p}=\|A\|_{p}$ for all quaternionic unitary $U,V$ of compatible sizes. In accordance with the standard terminology in the complex setting, we call $\|A\|_{1}$ the trace norm, $\|A\|_{2}$ the Frobenius norm, and $\|A\|_{\infty}$ the spectral norm.
By convention, from this point on, we write $\|A\|$ to denote spectral norm.
Proposition 2.8 (Symplectic embedding duplicates singular values).
Let $A\in\mathbb{H}^{m\times n}$. If $\sigma_{1}\geq\cdots\geq\sigma_{r}>0$ are the nonzero singular values of $A$, then the singular values of $\Phi(A)\in\mathbb{C}^{2m\times 2n}$ are
$\sigma_{1},\sigma_{1},\;\sigma_{2},\sigma_{2},\;\dots,\;\sigma_{r},\sigma_{r},$
followed by zeros to reach size $\min\{2m,2n\}$.
Corollary 2.9.
For $1\leq p\leq\infty$,
$\|A\|_{p}=2^{-1/p}\,\bigl\|\Phi(A)\bigr\|_{p}.$
2.4 Function theory preliminaries and notation
Let $D\subset\mathbb{R}^{k}$ be a real measurable set with Lebesgue measure $|D|$. We can consider $\mathbb{H}^{m\times n}\cong\mathbb{R}^{4mn}$ using Cartesian coordinates and say that a quaternion matrix function $f:D\to\mathbb{H}^{m\times n}$ is measurable if and only if all its real coordinates are measurable. Equivalently, using real-linearity and continuity of $\Phi$, $f:D\to\mathbb{H}^{m\times n}$ is measurable if and only if $\Phi\circ f:D\to\mathbb{C}^{2m\times 2n}$ is measurable: in the rest of the manuscript, for our purposes, the measurable domain $D$ is assumed to have finite and positive measure.
For asymptotic spectral and singular value theory of quaternion matrix sequences, we need to use the Schatten norms of measurable matrix-valued functions, their singular and canonical eigenvalue fields, and ensure that they are measurable. Fortunately, in analogy with the complex setting, analogous properties hold true.
Lemma 2.10.
Let $f:D\to\mathbb{H}^{r\times r}$ be a measurable function. Then for any $p\in[1,\infty]$, the Schatten norm map $x\mapsto\|f(x)\|_{p}$ is measurable; the singular value map $x\mapsto(\sigma_{1}(f(x)),\ldots,\sigma_{r}(f(x)))$ is measurable; and if $\lambda_{1}^{+}(x),\ldots,\lambda_{r}^{+}(x)$ denote the canonical eigenvalues of $f(x)$ (as in Theorem 2.5), then for every continuous symmetric $g:\mathbb{C}^{r}\to\mathbb{C}$, the composite $x\mapsto g(\lambda_{1}^{+}(x),\ldots,\lambda_{r}^{+}(x))$ is measurable.
Proof.
The measurability of $x\mapsto\Phi(f(x))$ is by definition. The singular value map in $\mathbb{C}^{k\times k}$ is continuous, hence measurable; the Schatten norms are continuous functions of singular values. For canonical eigenvalues, the coefficients of the characteristic polynomial of $\Phi(f(x))$ are measurable (polynomials of entries), and the multiset of roots depends Borel-measurably on the coefficients on bounded sets; selecting the $r$ roots with $\Im z\geq 0$ is a Borel operation on conjugate pairs. See [5]. ∎
Quaternion matrix-valued $L^{p}$ spaces are used to define Toeplitz matrices:
Definition 2.11 ($L^{p}$ spaces and norms).
Given $1\leq p<\infty$ and measurable $f:D\to\mathbb{H}^{m\times n}$, the $L^{p}$ norm of $f$ is
$\|f\|_{L^{p}(D;\mathbb{H}^{m\times n})}:=\Big(\int_{D}\|f(x)\|_{p}^{p}\,dx\Big)^{1/p},\qquad\|f\|_{L^{\infty}(D;\mathbb{H}^{m\times n})}:=\operatorname*{ess\,sup}_{x\in D}\|f(x)\|.$
Let $L^{p}(D;\mathbb{H}^{m\times n})$ be the space of measurable functions with finite $L^{p}$ norm modulo a.e.(almost everywhere) equality.
Remark 2.12 (Compatibility with $\Phi$).
By Corollary 2.9, for $1\leq p\leq\infty$,
$\|f\|_{L^{p}(D;\mathbb{H}^{m\times n})}=2^{-1/p}\,\|\Phi\circ f\|_{L^{p}(D;\mathbb{C}^{2m\times 2n})},$
with the usual convention $2^{1/\infty}=1$.
When dealing with spectral distributions of matrix sequences, we use the following set.
2.5 Left, right, and sandwich trigonometric polynomials
We recover the notion of trigonometric single-axis trigonometric polynomials and a few basic facts to be used in the subsequent derivations. Fix an imaginary unit $\mathbf{i}\in\operatorname{Im}\mathbb{H}$ and the slice $\mathbb{C}_{\mathbf{i}}$. We remark that from now on, by convention, we are going to define all quaternion exponential kernels in the slice $\mathbb{C}_{\mathbf{i}}$.
Consider the torus $\mathbb{T}^{d}=[-\pi,\pi)^{d}$ endowed with the Lebesgue measure $d\bm{\theta}$. For $\bm{m}=(m_{1},\ldots,m_{d})\in\mathbb{Z}^{d}$ and $\bm{\theta}=(\theta_{1},\ldots,\theta_{d})\in\mathbb{T}^{d}$, we set $\langle\bm{m},\bm{\theta}\rangle:=\sum_{j=1}^{d}m_{j}\theta_{j}$. Due to non-commutativity of the quaternion algebra, by first-principles, we define polynomials that might have the exponential monomial applied on the left, right, or both, to the coefficients. We call the latter case “sandwich” if we do not need to specify the exact disposition of variables.
More formally, given an ordered disjoint partition $S_{L}\dot{\cup}S_{R}=[d]$, define
$\langle\bm{m},\bm{\theta}\rangle_{S_{L}}:=\sum_{j\in S_{L}}m_{j}\theta_{j},\qquad\langle\bm{m},\bm{\theta}\rangle_{S_{R}}:=\sum_{j\in S_{R}}m_{j}\theta_{j},$
and we denote left/right exponential kernels by
$L_{S_{L}}(\bm{\theta},\bm{m}):=\exp\bigl(-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle_{S_{L}}\bigr),\qquad R_{S_{R}}(\bm{\theta},\bm{m}):=\exp\bigl(-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle_{S_{R}}\bigr).$
The purely left and purely right partitions are of special interest. The left case is defined $L_{[d]}(\bm{\theta},\bm{m})=\exp(-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle)$, $R_{\varnothing}\equiv 1$. Purely right is instead $L_{\varnothing}\equiv 1$, $R_{[d]}(\bm{\theta},\bm{m})=\exp(-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle)$.
The definition of trigonometric polynomials depends on the kernel choice, as described below.
Definition 2.13 (Left, right, and sandwich trigonometric polynomials).
Let $A_{\bm{m}}\in\mathbb{H}^{m\times n}$ and $\mathcal{F}\subset\mathbb{Z}^{d}$ be finite. Define
$p_{\mathrm{L}}(\bm{\theta})=\sum_{\bm{m}\in\mathcal{F}}e^{-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle}A_{\bm{m}},\qquad p_{\mathrm{R}}(\bm{\theta})=\sum_{\bm{m}\in\mathcal{F}}A_{\bm{m}}e^{-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle},$
and, for a fixed partition $S_{L}\dot{\cup}S_{R}=[d]$,
$p^{(S_{L},S_{R})}(\bm{\theta})=\sum_{\bm{m}\in\mathcal{F}}L_{S_{L}}(\bm{\theta},\bm{m})\,A_{\bm{m}}\,R_{S_{R}}(\bm{\theta},\bm{m}).$
We call $p^{(S_{L},S_{R})}$ a sandwich trigonometric polynomial. The degree is $\deg p:=\max_{\bm{m}\in\mathcal{F}}\|\bm{m}\|_{\infty}$.
The Fourier coefficients and partial sums mimic the complex case, taking into account the kernel positions.
Definition 2.14 (Sandwich Fourier coefficients and partial sums).
For $F\in L^{1}(\mathbb{T}^{d};\mathbb{H}^{m\times n})$ and $\bm{m}\in\mathbb{Z}^{d}$, define
$\widehat{F}^{(S_{L},S_{R})}(\bm{m}):=(2\pi)^{-d}\int_{\mathbb{T}^{d}}L_{S_{L}}(\bm{\theta},\bm{m})\,F(\bm{\theta})\,R_{S_{R}}(\bm{\theta},\bm{m})\,d\bm{\theta}\in\mathbb{H}^{m\times n}.$
For $N\in\mathbb{N}$, the rectangular partial sums are
$S_{N}^{(S_{L},S_{R})}F(\bm{\theta}):=\sum_{\|\bm{m}\|_{\infty}\leq N}L_{S_{L}}(\bm{\theta},\bm{m})\,\widehat{F}^{(S_{L},S_{R})}(\bm{m})\,R_{S_{R}}(\bm{\theta},\bm{m}).$
The left and right cases are obtained by $S_{L}=[d],S_{R}=\varnothing$ and $S_{L}=\varnothing,S_{R}=[d]$, respectively.
Clearly, the Fourier coefficients of trigonometric polynomials are the coefficients themselves and this is useful for notational purposes.
Lemma 2.15 (Transport under the symplectic embedding).
Write $F(\bm{\theta})=Z(\bm{\theta})+W(\bm{\theta})\mathbf{j}$ with $Z,W:\mathbb{T}^{d}\to\mathbb{C}_{\mathbf{i}}^{m\times n}$. Define the sign-flip $\tilde{\bm{m}}\in\mathbb{Z}^{d}$ by $\tilde{m}_{j}=m_{j}$ if $j\in S_{L}$ and $\tilde{m}_{j}=-m_{j}$ if $j\in S_{R}$. Then, for every $\bm{m}\in\mathbb{Z}^{d}$,
$\Phi\!\bigl(\widehat{F}^{(S_{L},S_{R})}(\bm{m})\bigr)=\begin{bmatrix}\widehat{Z}(\bm{m})&\widehat{W}(\tilde{\bm{m}})\\[2.0pt] -\widehat{\overline{W}}(-\tilde{\bm{m}})&\widehat{\overline{Z}}(-\bm{m})\end{bmatrix}$,
where the hats on the right denote ordinary complex matrix Fourier coefficients on $\mathbb{T}^{d}$.
Proof.
Use $\mathbf{j}\alpha=\overline{\alpha}\,\mathbf{j}$ for all $\alpha\in\mathbb{C}_{\mathbf{i}}$ and the identities $\mathbf{j}e^{-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle}=e^{+\mathbf{i}\langle\bm{m},\bm{\theta}\rangle}\mathbf{j}$, $e^{-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle}\mathbf{j}=\mathbf{j}e^{+\mathbf{i}\langle\bm{m},\bm{\theta}\rangle}$, to rewrite
$L_{S_{L}}(\bm{\theta},\bm{m})\,F(\bm{\theta})\,R_{S_{R}}(\bm{\theta},\bm{m})=Z(\bm{\theta})\,e^{-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle}+W(\bm{\theta})\,e^{-\mathbf{i}\langle\tilde{\bm{m}},\bm{\theta}\rangle}\,\mathbf{j}.$
Then apply $\Phi$ and integrate entrywise. ∎
In addition, as it can be easily checked via the embedding trick, the following result holds.
Proposition 2.16 (Density in $L^{p}$).
For $1\leq p<\infty$, the linear span of sandwich trigonometric polynomials $\{p^{(S_{L},S_{R})}\}$ is dense in $L^{p}(\mathbb{T}^{d};\mathbb{H}^{m\times n})$.
In the single-axis setting, using (2.2) to switch from different trigonometric classes is standard, and rectangular matrix-valued polynomials follow analogous arguments. In the signal analysis setting, this can be viewed as a special and trivial case of the splitting technique of general discrete Fourier transforms with two axes (see [23]), to convert sandwich kernels to a unidirectional one, such as the classical right kernel. For matrix-valued polynomials and Fourier coefficients, we indeed have
Proposition 2.17 (Reduction to the right case).
Let $S_{L}\dot{\cup}S_{R}=[d]$ and define
$\tilde{m}_{j}=\begin{cases}m_{j},&j\in S_{L},\\ -m_{j},&j\in S_{R}.\end{cases}$
For finite $\mathcal{F}\subset\mathbb{Z}^{d}$ and $A_{\bm{m}}\in\mathbb{H}^{m\times n}$, write $A_{\bm{m}}=Z_{\bm{m}}+W_{\bm{m}}\mathbf{j}$ with $Z_{\bm{m}},W_{\bm{m}}\in\mathbb{C}_{\mathbf{i}}^{m\times n}$. With $L_{S_{L}},R_{S_{R}}$ as in Definitions 2.13, 2.14, the sandwich polynomial
$p^{(S_{L},S_{R})}(\bm{\theta})=\sum_{\bm{m}\in\mathcal{F}}L_{S_{L}}(\bm{\theta},\bm{m})\,A_{\bm{m}}\,R_{S_{R}}(\bm{\theta},\bm{m})$
reduces to a sum of two right polynomials (in the slice $\mathbb{C}_{\mathbf{i}}$):
$p^{(S_{L},S_{R})}(\bm{\theta})=\sum_{\bm{m}\in\mathcal{F}}\Bigl(Z_{\bm{m}}e^{-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle}+W_{\bm{m}}e^{-\mathbf{i}\langle\tilde{\bm{m}},\bm{\theta}\rangle}\mathbf{j}\Bigr).$ (2.6)
Moreover, for $F=Z+W\mathbf{j}\in L^{1}(\mathbb{T}^{d};\mathbb{H}^{m\times n})$ with $Z,W:\mathbb{T}^{d}\to\mathbb{C}_{\mathbf{i}}^{m\times n}$, the sandwich Fourier coefficients
$\widehat{F}^{(S_{L},S_{R})}(\bm{m})=(2\pi)^{-d}\int_{\mathbb{T}^{d}}L_{S_{L}}(\bm{\theta},\bm{m})\,F(\bm{\theta})\,R_{S_{R}}(\bm{\theta},\bm{m})\,d\bm{\theta}$
satisfy
$\widehat{F}^{(S_{L},S_{R})}(\bm{m})=\widehat{Z}(\bm{m})+\widehat{W}(\tilde{\bm{m}})\,\mathbf{j},$ (2.7)
where the hats on the right are the usual complex (right) Fourier coefficients on $\mathbb{T}^{d}$ in the slice $\mathbb{C}_{\mathbf{i}}$. In particular,
$\widehat{F}_{\mathrm{L}}(\bm{m})=\overline{\,\widehat{\,\overline{F}\,}_{\mathrm{R}}(-\bm{m})\,}.$ (2.8)
Proof.
Fix $\bm{m}\in\mathbb{Z}^{d}$ and write $A_{\bm{m}}=Z_{\bm{m}}+W_{\bm{m}}\mathbf{j}$ with $Z_{\bm{m}},W_{\bm{m}}\in\mathbb{C}_{\mathbf{i}}^{m\times n}$. Set $a=\langle\bm{m},\bm{\theta}\rangle_{S_{L}}$ and $b=\langle\bm{m},\bm{\theta}\rangle_{S_{R}}$. Since $e^{-\mathbf{i}a},e^{-\mathbf{i}b}\in\mathbb{C}_{\mathbf{i}}$ commute with $Z_{\bm{m}},W_{\bm{m}}$ and $\mathbf{j}\alpha=\overline{\alpha}\,\mathbf{j}$ for all $\alpha\in\mathbb{C}_{\mathbf{i}}$, we compute
$L_{S_{L}}A_{\bm{m}}R_{S_{R}}=e^{-\mathbf{i}a}(Z_{\bm{m}}+W_{\bm{m}}\mathbf{j})e^{-\mathbf{i}b}=Z_{\bm{m}}e^{-\mathbf{i}(a+b)}+W_{\bm{m}}\,(e^{-\mathbf{i}a}\mathbf{j})e^{-\mathbf{i}b}=Z_{\bm{m}}e^{-\mathbf{i}(a+b)}+W_{\bm{m}}\,e^{-\mathbf{i}(a-b)}\mathbf{j}.$
Because $a+b=\langle\bm{m},\bm{\theta}\rangle$ and $a-b=\langle\tilde{\bm{m}},\bm{\theta}\rangle$, this yields (2.6).
Now let $F=Z+W\mathbf{j}\in L^{1}(\mathbb{T}^{d};\mathbb{H}^{m\times n})$ with $Z,W:\mathbb{T}^{d}\to\mathbb{C}_{\mathbf{i}}^{m\times n}$. Since the decomposition $F\mapsto(Z,W)$ is given by a bounded linear map $\mathbb{H}^{m\times n}\to\mathbb{C}_{\mathbf{i}}^{m\times n}\times\mathbb{C}_{\mathbf{i}}^{m\times n}$, we infer $Z,W\in L^{1}$. Applying the previous algebra pointwise and integrating entrywise,
$\widehat{F}^{(S_{L},S_{R})}(\bm{m})=\frac{1}{(2\pi)^{d}}\!\int_{\mathbb{T}^{d}}\bigl(Z(\bm{\theta})e^{-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle}+W(\bm{\theta})e^{-\mathbf{i}\langle\tilde{\bm{m}},\bm{\theta}\rangle}\mathbf{j}\bigr)\,d\bm{\theta}=\widehat{Z}(\bm{m})+\widehat{W}(\tilde{\bm{m}})\,\mathbf{j},$
which is (2.7).
For (2.8), by definition $\widehat{F}_{\mathrm{L}}(\bm{m})=\frac{1}{(2\pi)^{d}}\!\int_{\mathbb{T}^{d}}e^{-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle}F(\bm{\theta})\,d\bm{\theta}$. Taking quaternionic conjugates, using $\overline{AB}=\overline{B}\,\overline{A}$ and $\overline{e^{-\mathbf{i}t}}=e^{\mathbf{i}t}$,
$\overline{\widehat{F}_{\mathrm{L}}(\bm{m})}=\frac{1}{(2\pi)^{d}}\!\int_{\mathbb{T}^{d}}\overline{F(\bm{\theta})}\,e^{\mathbf{i}\langle\bm{m},\bm{\theta}\rangle}\,d\bm{\theta}=\frac{1}{(2\pi)^{d}}\!\int_{\mathbb{T}^{d}}\overline{F(\bm{\theta})}\,e^{-\mathbf{i}\langle-\bm{m},\bm{\theta}\rangle}\,d\bm{\theta}=\widehat{\,\overline{F}\,}_{\mathrm{R}}(-\bm{m}),$
and conjugating both sides gives the claim. ∎
Remark 2.18.
The Proposition applied to the purely left kernel gives us $\widehat{F}_{\mathrm{R}}(\bm{m})=\overline{\,\widehat{\,\overline{F}\,}_{\mathrm{L}}(-\bm{m})\,}$.
3 Asymptotic spectral distribution of quaternion matrix sequences
In this section we present the asymptotic spectral and singular value distribution formalism for quaternion matrix sequences. We adopt the point of view of Theorem 2.5 and Proposition 2.8. Regarding the spectral analysis, canonical eigenvalues encode the eigenvalue information of quaternion matrices, while, for singular values, the analysis reduces to the complex case under the symplectic embedding $\Phi$. This allows us to inherit the well-developed framework for asymptotic spectral distributions of complex matrix sequences as developed in [5, 9, 21, 54, 62] and references therein. Throughout, eigenvalues are listed with algebraic multiplicity and singular values are listed in nonincreasing order. Measurability of the spectral mappings $x\mapsto\lambda_{i}(G(x))$ and $x\mapsto\sigma_{i}(H(x))$ needed below follows from Lemma 2.10.
Eigenvalue distributions via the symplectic embedding.
Definition 3.1 (Embedding-based spectral symbol).
Let $A_{n}\in\mathbb{H}^{d_{n}\times d_{n}}$ be square quaternionic matrices with $d_{n}\to\infty$. Let $D\subset\mathbb{R}^{k}$ be measurable with $0<|D|<\infty$, and let $G:D\to\mathbb{C}^{q\times q}$ be a measurable matrix-valued function. We say that the quaternionic sequence $\{A_{n}\}_{n}$ is distributed in the sense of the eigenvalues as $G$ on $D$, and we write
$\{A_{n}\}_{n}\sim_{\lambda}(G,D),$
if and only if the complex embeddings satisfy $\{\Phi(A_{n})\,\}_{n}\sim_{\lambda}(G,D)$ in the classical sense. Concretely, for every $F\in C_{c}(\mathbb{C})$,
$\lim_{n\to\infty}\frac{1}{2d_{n}}\sum_{j=1}^{2d_{n}}F\!\bigl(\lambda_{j}(\Phi(A_{n}))\bigr)\;=\;\frac{1}{|D|}\int_{D}\frac{1}{q}\sum_{i=1}^{q}F\!\bigl(\lambda_{i}(G(x))\bigr)\,dx,$ (3.1)
where the eigenvalues $\lambda_{j}(\Phi(A_{n}))$ and $\lambda_{i}(G(x))$ are listed with algebraic multiplicity.
Let $\lambda_{1}^{+}(A_{n}),\dots,\lambda_{d_{n}}^{+}(A_{n})$ denote the canonical eigenvalues of $A_{n}$ (the upper half-plane selector; see Theorem 2.5). Then the eigenvalues of $\Phi(A_{n})$ occur in conjugate pairs $\{z,\overline{z}\}$; real eigenvalues appear as pairs $\{z,z\}$. Therefore, for every $F\in C_{c}(\mathbb{C})$,
$\frac{1}{2d_{n}}\sum_{j=1}^{2d_{n}}F\bigl(\lambda_{j}(\Phi(A_{n}))\bigr)\;=\;\frac{1}{d_{n}}\sum_{j=1}^{d_{n}}\widetilde{F}\bigl(\lambda^{+}_{j}(A_{n})\bigr),\qquad\widetilde{F}(z):=\tfrac{1}{2}\bigl(F(z)+F(\overline{z})\bigr).$ (3.2)
Equation (3.2) shows that the embedding-based definition averages each conjugate pair $\{z,\overline{z}\}$ via $\widetilde{F}$. In particular, the two sides coincide when $F$ is conjugation-symmetric, i.e., $F(\overline{z})=F(z)$.
Alternative quaternion-side formulation for eigenvalues. If an entirely quaternionic formulation is preferred, then the following definition can be taken into consideration.
Definition 3.2 (Canonical-eigenvalue distribution).
Let $A_{n}\in\mathbb{H}^{d_{n}\times d_{n}}$ with $d_{n}\to\infty$, and let $\lambda_{1}^{+}(A_{n}),\dots,\lambda_{d_{n}}^{+}(A_{n})$ be the canonical eigenvalues as in Theorem 2.5. Let $D\subset\mathbb{R}^{k}$ be measurable with $0<|D|<\infty$, and let $G\in L^{0}(D;\mathbb{C}^{q\times q})$. We write that $\{A_{n}\}\sim_{\lambda^{+}}(G,D)$ if, for every $F\in C_{c}(\mathbb{C})$ with $F(\overline{z})=F(z)$,
$\lim_{n\to\infty}\frac{1}{d_{n}}\sum_{j=1}^{d_{n}}F\bigl(\lambda_{j}^{+}(A_{n})\bigr)\;=\;\frac{1}{|D|}\int_{D}\frac{1}{q}\sum_{i=1}^{q}F\bigl(\lambda_{i}(G(x))\bigr)\,dx.$
Remark 3.3.
By (3.2), if $\{A_{n}\}\sim_{\lambda}(G,D)$ in the sense of Definition 3.1, then $\{A_{n}\}\sim_{\lambda^{+}}(G,D)$ in the sense of Definition 3.2. Conversely, if for a.e. $x\in D$ the multiset $\{\lambda_{i}(G(x))\}_{i=1}^{q}$ is invariant under complex conjugation, then $\{A_{n}\}\sim_{\lambda^{+}}(G,D)$ implies $\{A_{n}\}\sim_{\lambda}(G,D)$.
Remark 3.4.
The same statement holds verbatim for the embedding-based notion $\sim_{\lambda}$ of Definition 3.1 (no conjugation-symmetry restriction on $F$). Thus the eigenvalue symbol is determined only up to $D$-measure-preserving reparametrizations of the spatial variable.
Singular value distributions via the symplectic embedding.
Definition 3.5 (Embedding-based singular value symbol).
Let $A_{n}\in\mathbb{H}^{m_{n}\times n_{n}}$ be rectangular quaternionic matrices and set $r_{n}:=\min\{m_{n},n_{n}\}\to\infty$. Let $D\subset\mathbb{R}^{k}$ be measurable with $0<|D|<\infty$, and let $H\in L^{0}(D;\mathbb{C}^{s\times t})$ be measurable. We write that $\{A_{n}\}_{n}$ is distributed in singular values as $H$ on $D$, and we write
$\{A_{n}\}_{n}\sim_{\sigma}(H,D),$
if and only if $\{\,\Phi(A_{n})\,\}_{n}\sim_{\sigma}(H,D)$ in the classical sense, i.e., for every $\psi\in C_{c}([0,\infty))$,
$\lim_{n\to\infty}\frac{1}{2r_{n}}\sum_{i=1}^{2r_{n}}\psi\!\bigl(\sigma_{i}(\Phi(A_{n}))\bigr)\;=\;\frac{1}{|D|}\int_{D}\frac{1}{\min\{s,t\}}\sum_{i=1}^{\min\{s,t\}}\psi\!\bigl(\sigma_{i}(H(x))\bigr)\,dx.$ (3.3)
Remark 3.6 (Duplication of singular values under the embedding).
By Proposition 2.8, the singular values of $\Phi(A_{n})$ are exactly those of $A_{n}$, each duplicated. Consequently, one can evaluate the empirical singular value distributions of quaternionic matrices directly on the quaternionic side without passing to the complex representation; see (3.4) below.
Normalization identities. The following equalities make the scaling in (3.3) explicit.
$\frac{1}{2r_{n}}\sum_{i=1}^{2r_{n}}\psi\bigl(\sigma_{i}(\Phi(A_{n}))\bigr)\;=\;\frac{1}{r_{n}}\sum_{i=1}^{r_{n}}\psi\bigl(\sigma_{i}(A_{n})\bigr),\qquad\forall\,\psi\in C_{c}([0,\infty)).$ (3.4)
Identity (3.4) is an immediate consequence of Proposition 2.8.
For the properties of a.c.s. we need the following definition
Definition 3.7 (Sparsely unbounded quaternion matrix sequences).
Let $A_{n}\in\mathbb{H}^{m_{n}\times n_{n}}$ be (rectangular) quaternionic matrices and set $r_{n}:=\min\{m_{n},n_{n}\}\to\infty$. Denote the singular values of $A_{n}$ (by definition, the square roots of the eigenvalues of $A_{n}^{*}A_{n}$, with the quaternionic conjugate transpose) in nonincreasing order by $\sigma_{1}(A_{n})\geq\cdots\geq\sigma_{r_{n}}(A_{n})\geq 0$. We write that the sequence $\{A_{n}\}_{n}$ is sparsely unbounded (s.u.) if there exists a function $r:(0,\infty)\to[0,1]$ with $\lim_{M\to\infty}r(M)=0$ such that, for every $M>0$, there is $n_{M}$ with the property that, for all $n\geq n_{M}$,
$\frac{1}{r_{n}}\,\#\Bigl\{\,i\in\{1,\dots,r_{n}\}:\ \sigma_{i}(A_{n})>M\,\Bigr\}\;\leq\;r(M).$
Here $#(\cdot)$ denotes the cardinality.
3.1 Definition of a.c.s.
We introduce the notion of rectangular a.c.s. over $\mathbb{H}$. Throughout this section, $\|\cdot\|$ denotes the operator (spectral) norm on quaternionic matrices (cf. Definition 2.7 for Schatten $p$-norms).
Definition 3.8 (Approximating class of sequences).
Let $\{A_{n}\}_{n}$ be a matrix sequence with $A_{n}\in\mathbb{H}^{d_{n}\times e_{n}}$ and set $r_{n}:=d_{n}\wedge e_{n}\to\infty$. A class of matrix sequences $\{\{B_{n,m}\}_{n}:\ m\}$ with $B_{n,m}\in\mathbb{H}^{d_{n}\times e_{n}}$ is an approximating class of sequences for $\{A_{n}\}_{n}$ if, for every fixed $m$, there exist $R_{n,m},N_{n,m}\in\mathbb{H}^{d_{n}\times e_{n}}$ and $n_{m}$ such that, for all $n\geq n_{m}$,
$A_{n}$ $=B_{n,m}+R_{n,m}+N_{n,m},$ (3.5)
$\frac{\operatorname{rank}_{\mathbb{H}}(R_{n,m})}{r_{n}}$ $\leq c(m),\qquad\|N_{n,m}\|\leq\omega(m),$ (3.6)
where $c(m)\to 0$ and $\omega(m)\to 0$ as $m\to\infty$. We write $\{B_{n,m}\}_{n}\xrightarrow[\mathbb{H}]{\mathrm{a.c.s.}}\{A_{n}\}_{n}$ when the quaternion side must be explicit, otherwise $\{B_{n,m}\}_{n}\xrightarrow{\mathrm{a.c.s.}}\{A_{n}\}_{n}$, and it must be emphasized that the limit is taken with respect to the external variable $m$.
For later use we record the basic norm/rank relations induced by the symplectic embedding.
Lemma 3.9 (Norm and rank under $\Phi$).
For all $X\in\mathbb{H}^{d\times e}$ and $1\leq p\leq\infty$,
$\|\Phi(X)\|=\|X\|,\qquad\|\Phi(X)\|_{S_{p}}=2^{1/p}\|X\|_{p},\qquad\operatorname{rank}_{\mathbb{C}}\!\bigl(\Phi(X)\bigr)=2\,\operatorname{rank}_{\mathbb{H}}(X).$
Proof. The singular values of $\Phi(X)$ are those of $X$, each duplicated (Proposition 2.8); the equalities follow, using also Corollary 2.9 and $\|\cdot\|=\|\cdot\|_{\infty}$. ∎
Since $\Phi$ preserves the operator norm and duplicates nonzero singular values, the following equivalence with the complex a.c.s. framework (see [50]) holds.
Theorem 3.10 (Embedding equivalence).
Let $\{A_{n}\}_{n}\subset\mathbb{H}^{d_{n}\times e_{n}}$ with $r_{n}:=d_{n}\wedge e_{n}\to\infty$. Use the rank normalizations $\operatorname{rank}_{\mathbb{H}}(\cdot)/r_{n}$ on the quaternionic side and $\operatorname{rank}_{\mathbb{C}}(\cdot)/(2r_{n})$ on the embedded complex side. Then the following are equivalent:
  1. (a)
    $\{A_{n}\}_{n}$ admits a quaternionic a.c.s. $\{\{B_{n,m}\}_{n}:\ m\}$ as in Definition 3.8.
  2. (b)
    The double-size matrix sequence $\{\Phi(A_{n})\}_{n}$ admits a complex a.c.s. $\{\{C_{n,m}\}_{n}:\ m\}$ (as in [50]) with the same $n_{m}$ and control functions, measured with the above normalization.
[…]
Proof.
We use Lemma 3.9.
[…]
$C_{n,m}:=\Phi(B_{n,m}),\quad S_{n,m}:=\Phi(R_{n,m}),\quad T_{n,m}:=\Phi(N_{n,m}).$
Then
$\Phi(A_{n})=C_{n,m}+S_{n,m}+T_{n,m},\qquad\frac{\operatorname{rank}_{\mathbb{C}}(S_{n,m})}{2r_{n}}=\frac{2\,\operatorname{rank}_{\mathbb{H}}(R_{n,m})}{2r_{n}}\leq c(m),\quad\|T_{n,m}\|=\|N_{n,m}\|\leq\omega(m),$
so $\{C_{n,m}\}_{n,m}$ is a complex rectangular a.c.s. with the same $n_{m}$, $c(m)$, and $\omega(m)$.
[…]
$\mathsf{J}_{k}:=\begin{bmatrix}0&I_{k}\\[-1.0pt] -I_{k}&0\end{bmatrix},\qquad\tau_{d_{n},e_{n}}(X):=-\,\mathsf{J}_{d_{n}}\,\overline{X}\,\mathsf{J}_{e_{n}},\qquad\Pi_{d_{n},e_{n}}:=\tfrac{1}{2}(\mathrm{id}+\tau_{d_{n},e_{n}}).$
Then $\tau_{d_{n},e_{n}}$ is a conjugate-linear isometric involution with fixed-point space $\mathrm{Fix}(\tau_{d_{n},e_{n}})=\mathrm{Ran}(\Phi)$; hence $\Pi_{d_{n},e_{n}}$ is the (real-linear) projector onto $\mathrm{Ran}(\Phi)$ and satisfies $\|\Pi_{d_{n},e_{n}}(X)\|\leq\|X\|$. For $n\geq n_{m}$ assume that
$\Phi(A_{n})=C_{n,m}+S_{n,m}+T_{n,m},\qquad\frac{\operatorname{rank}_{\mathbb{C}}(S_{n,m})}{2r_{n}}\leq c(m),\ \ \|T_{n,m}\|\leq\omega(m).$ (3.7)
Project (3.7) onto $\mathrm{Ran}(\Phi)$:
$\Phi(A_{n})=\Pi_{d_{n},e_{n}}(C_{n,m})+\Pi_{d_{n},e_{n}}(S_{n,m})+\Pi_{d_{n},e_{n}}(T_{n,m}),$
with
$\|\Pi_{d_{n},e_{n}}(T_{n,m})\|\leq\omega(m),\qquad\operatorname{rank}_{\mathbb{C}}\!\bigl(\Pi_{d_{n},e_{n}}(S_{n,m})\bigr)\leq\operatorname{rank}_{\mathbb{C}}(S_{n,m})+\operatorname{rank}_{\mathbb{C}}\!\bigl(\tau_{d_{n},e_{n}}(S_{n,m})\bigr)\leq 2\,\operatorname{rank}_{\mathbb{C}}(S_{n,m}).$
Since each term lies in $\mathrm{Ran}(\Phi)$, pull back via the inverse of $\Phi$ on its range:
$B_{n,m}:=\Phi^{-1}\!\bigl(\Pi_{d_{n},e_{n}}(C_{n,m})\bigr),\quad R_{n,m}:=\Phi^{-1}\!\bigl(\Pi_{d_{n},e_{n}}(S_{n,m})\bigr),\quad N_{n,m}:=\Phi^{-1}\!\bigl(\Pi_{d_{n},e_{n}}(T_{n,m})\bigr).$
Then $A_{n}=B_{n,m}+R_{n,m}+N_{n,m}$ and
$\frac{\operatorname{rank}_{\mathbb{H}}(R_{n,m})}{r_{n}}=\frac{\operatorname{rank}_{\mathbb{C}}(\Pi_{d_{n},e_{n}}(S_{n,m}))}{2r_{n}}\leq 2\,c(m),\qquad\|N_{n,m}\|=\|\Pi_{d_{n},e_{n}}(T_{n,m})\|\leq\omega(m).$
Thus $\{\{B_{n,m}\}_{n}:\ m\}$ is a quaternionic a.c.s. with the same $n_{m}$, the same $\omega(m)$, and rank control at most $2c(m)$. ∎
Consequences. All standard a.c.s. properties transfer verbatim with $r_{n}=d_{n}\wedge e_{n}$:
  1. ACSH1.
    $\{A_{n}\}\sim_{\sigma}F$ if and only if there exist matrix sequences $\{B_{n,m}\}_{n}\sim_{\sigma}F_{m}$ with $\{B_{n,m}\}_{n}\xrightarrow[\mathbb{H}]{\mathrm{a.c.s.}}\{A_{n}\}_{n}$ and $F_{m}\to F$ in measure.
  2. ACSH2.
    If every $A_{n}$ is Hermitian and square, then $\{A_{n}\}\sim_{\lambda}f$ if and only if there exist Hermitian sequences $\{B_{n,m}\}_{n}\sim_{\lambda}f_{m}$ with $\{B_{n,m}\}_{n}\xrightarrow[\mathbb{H}]{\mathrm{a.c.s.}}\{A_{n}\}_{n}$ and $f_{m}\to f$ in measure.
  3. ACSH3.
    Let $A_{n}\in\mathbb{H}^{d_{n}\times e_{n}}$, $A^{\prime}_{n}\in\mathbb{H}^{e_{n}\times f_{n}}$ and analogously $B_{n,m},B^{\prime}_{n,m}$. If $\{B_{n,m}\}_{n}\xrightarrow[\mathbb{H}]{\mathrm{a.c.s.}}\{A_{n}\}_{n}$ and $\{B^{\prime}_{n,m}\}_{n}\xrightarrow[\mathbb{H}]{\mathrm{a.c.s.}}\{A^{\prime}_{n}\}_{n}$, then
    $\{B_{n,m}^{*}\}_{n}\xrightarrow[\mathbb{H}]{\mathrm{a.c.s.}}\{A_{n}^{*}\}_{n},$
    $\{\alpha B_{n,m}+\beta B^{\prime}_{n,m}\}_{n}\xrightarrow[\mathbb{H}]{\mathrm{a.c.s.}}\{\alpha A_{n}+\beta A^{\prime}_{n}\}_{n}\quad\text{(for left scalars $\alpha,\beta\in\mathbb{H}$),}$
    $\{B_{n,m}B^{\prime}_{n,m}\}_{n}\xrightarrow[\mathbb{H}]{\mathrm{a.c.s.}}\{A_{n}A^{\prime}_{n}\}_{n}\quad\text{(whenever $\{A_{n}\}_{n},\{A^{\prime}_{n}\}_{n}$ are sparsely unbounded,}$
    $\{B_{n,m}C_{n}\}_{n}\xrightarrow[\mathbb{H}]{\mathrm{a.c.s.}}\{A_{n}C_{n}\}_{n}\quad\text{if $\{C_{n}\}_{n}$ is sparsely unbounded.}$
  4. ACSH4.
    If $A_{n}=[A_{n,ij}]_{i,j=1}^{s}$ and $B_{n}^{(m)}=[B^{(m)}_{n,ij}]_{i,j=1}^{s}$ with $\{B^{(m)}_{n,ij}\}_{n}\xrightarrow[\mathbb{H}]{\mathrm{a.c.s.}}\{A_{n,ij}\}_{n}$ for every $i,j$, then $\{B_{n}^{(m)}\}_{n}\xrightarrow[\mathbb{H}]{\mathrm{a.c.s.}}\{A_{n}\}_{n}$.
  5. ACSH5.
    (Practical norm test.) Let $p\in[1,\infty]$ and assume that for each $m$ there exists $n_{m}$ such that, for all $n\geq n_{m}$,
    $\|A_{n}-B_{n,m}\|_{p}\ \leq\ \varepsilon(m,n)\,r_{n}^{1/p},$
    with the convention $r_{n}^{1/\infty}=1$, where $\|\cdot\|_{p}$ denotes the Schatten $p$-norm (Definition 2.7), $\lim_{m\to\infty}\limsup_{n\to\infty}\varepsilon(m,n)=0$, and $r_{n}=d_{n}\wedge e_{n}$. Then $\{B_{n,m}\}_{n}\xrightarrow[\mathbb{H}]{\mathrm{a.c.s.}}\{A_{n}\}_{n}$.
4 Single-axis quaternion Toeplitz sequences
This section contains the basic asymptotic theory for multilevel block Toeplitz sequences generated by multivariate matrix-valued quaternion functions in $L^{1}$. Our starting point is, for most results, [35] for the case of unilevel scalar Toeplitz matrices generated by bounded functions. Unlike the complex case, the Fourier quaternion kernel is not unique and can depend on two orthogonal imaginary axes (see [17] for a discussion of the topic). Here, we focus our attention on a single-axis and consider single-axis Toeplitz sequences. Under the single-axis hypothesis, the question of how to apply the Fourier kernel remains due to the non-commutativity of quaternions; however, we will see that even if, numerically, a given generating function yields different quaternion Toeplitz matrices depending on the kernel, algebraically, every Toeplitz matrix sequence belongs to the same class, and this is independent of the construction of Fourier coefficients. Other than this remark, in this section we recover all the fundamental classical properties of Toeplitz sequences that one can find in standard GLT theory surveys, such as [20, Chapter 6], [21, Chapter 4], [5].
We write $\mathbb{T}^{d}=[-\pi,\pi)^{d}$ and use the Lebesgue measure $d\bm{\theta}$ on $\mathbb{T}^{d}$; all Fourier formulas below include the prefactor $(2\pi)^{-d}$.
4.1 $d$-level quaternion Toeplitz matrices
To construct quaternion Toeplitz matrices we use the kernel $L_{S_{L}}$ and $R_{S_{R}}$, applied respectively on the left and right, given in Definition 2.13, and the Fourier coefficients $F^{(S_{L},S_{R})}(\bm{m})$ are as in Definition 2.14.
Thus, for $\bm{n}=(n_{1},\dots,n_{d})\in\mathbb{N}^{d}$ and a fixed multi-index set
$\Lambda_{\bm{n}}:=\{0,\dots,n_{1}-1\}\times\cdots\times\{0,\dots,n_{d}-1\},$ (4.1)
endowed with the standard lexicographical order, we define the $d$-level (block) Toeplitz matrix as
$T_{\bm{n}}^{(S_{L},S_{R})}\!\bigl(F\bigr):=\bigl[\widehat{F}^{(S_{L},S_{R})}(\alpha-\beta)\bigr]_{\alpha,\beta\in\Lambda_{\mathbf{n}}}\in\mathbb{H}^{(N_{\bm{n}}s)\times(N_{\bm{n}}t)},\qquad N_{\bm{n}}=\prod_{\ell=1}^{d}n_{\ell}.$ (4.2)
The structure is analogous to that of its complex counterpart; however, the algebra this structure generates is different. First, we observe that all Toeplitz matrices belong to the same class. This is because the underlying kernel for Toeplitz matrices is interchangeable.
Proposition 4.1 (Right-kernel reduction for Toeplitz matrices).
Let $F=Z+W\mathbf{j}\in L^{1}(\mathbb{T}^{d};\mathbb{H}^{s\times t})$ with $Z,W:\mathbb{T}^{d}\to\mathbb{C}_{\mathbf{i}}^{s\times t}$ and fix an ordered partition $S_{L}\dot{\cup}S_{R}=[d]$. Consider
$(r_{S_{L}}(\bm{\theta}))_{j}=\begin{cases}-\theta_{j},&j\in S_{L},\\ \ \theta_{j},&j\in S_{R},\end{cases}$
and define
$H(\bm{\theta}):=Z(\bm{\theta})+\bigl(W\circ r_{S_{L}}\bigr)(\bm{\theta})\,\mathbf{j}.$ (4.3)
Then, for every $\bm{n}\in\mathbb{N}^{d}$,
$T_{\bm{n}}^{(S_{L},S_{R})}(F)=T_{\bm{n}}^{(R)}(H).$ (4.4)
In particular, for the left kernel ($S_{L}=[d],S_{R}=\varnothing$) case we get
$T_{\bm{n}}^{(L)}(F)=T_{\bm{n}}^{(R)}\!\Bigl(Z+\bigl(W\circ(-\mathrm{id})\bigr)\mathbf{j}\Bigr).$ (4.5)
Moreover, for a fixed $\bm{n}\in\mathbb{N}^{d}$ and the Cartesian set
$\Delta(\bm{n}):=\{\,\bm{k}\in\mathbb{Z}^{d}:\ -n_{j}+1\leq k_{j}\leq n_{j}-1,\ j\in[d]\,\},$
we have
$T_{\bm{n}}^{(L)}(F)=T_{\bm{n}}^{(R)}(F)\ \Longleftrightarrow\ \widehat{W}(\bm{k})=\widehat{W}(-\bm{k})\quad\forall\,\bm{k}\in\Delta(\bm{n}).$ (4.6)
In particular, at the matrix sequence level
$\{\,T_{\bm{n}}^{(L)}(F)\,\}_{\bm{n}}=\{\,T_{\bm{n}}^{(R)}(F)\,\}_{\bm{n}}\ \Longleftrightarrow\ \widehat{W}(\bm{k})=\widehat{W}(-\bm{k})\quad\forall\,\bm{k}\in\mathbb{Z}^{d},$ (4.7)
and the latter is equivalent to $W(\bm{\theta})=W(-\bm{\theta})$ a.e. on $\mathbb{T}^{d}$.
Proof.
Fix $\bm{k}\in\mathbb{Z}^{d}$. By equation (2.7), we write
$\widehat{F}^{(S_{L},S_{R})}(\bm{k})=\widehat{Z}(\bm{k})+\widehat{W}(\tilde{\bm{k}})\,\mathbf{j},$
where $\tilde{k}_{j}=k_{j}$ for $j\in S_{L}$ and $\tilde{k}_{j}=-k_{j}$ for $j\in S_{R}$. Given $H=Z+(W\circ r_{S_{L}})\mathbf{j}$, applying (2.7) with the right kernel ($S_{L}=\varnothing,S_{R}=[d]$) yields
$\widehat{H}^{(R)}(\bm{k})=\widehat{Z}(\bm{k})+\widehat{\,W\circ r_{S_{L}}\,}(-\bm{k})\,\mathbf{j},$
and since $\langle-\bm{k},r_{S_{L}}(\bm{\theta})\rangle=\langle\tilde{\bm{k}},\bm{\theta}\rangle$ and Lebesgue measure on $\mathbb{T}^{d}$ is invariant under $r_{S_{L}}$,
[…]
Thus $\widehat{H}^{(R)}(\bm{k})=\widehat{F}^{(S_{L},S_{R})}(\bm{k})$ for all $\bm{k}$, which implies
$\bigl[T_{\bm{n}}^{(S_{L},S_{R})}(F)\bigr]_{\alpha,\beta}=\widehat{F}^{(S_{L},S_{R})}(\alpha-\beta)=\widehat{H}^{(R)}(\alpha-\beta)=\bigl[T_{\bm{n}}^{(R)}(H)\bigr]_{\alpha,\beta},$
proving (4.4). Taking $S_{L}=[d]$ gives (4.5).
For the fixed-size criterion, $T_{\bm{n}}^{(L)}(F)=T_{\bm{n}}^{(R)}(F)$ iff $\widehat{F}^{(L)}(\bm{k})=\widehat{F}^{(R)}(\bm{k})$ for all difference indices $\bm{k}=\alpha-\beta$ that occur in $T_{\bm{n}}$, i.e., for all $\bm{k}\in\Delta(\bm{n})$; namely
$\widehat{Z}(\bm{k})+\widehat{W}(\bm{k})\mathbf{j}=\widehat{Z}(\bm{k})+\widehat{W}(-\bm{k})\mathbf{j}\quad\text{for all }\bm{k}\in\Delta(\bm{n}),$
which is equivalent to (4.6). At the matrix sequence level, the equality for all $\bm{n}$ holds if and only if the above identity holds for all $\bm{k}\in\mathbb{Z}^{d}$, yielding (4.7), which is equivalent to $W(\bm{\theta})=W(-\bm{\theta})$ a.e. ∎
4.2 Linearity, Adjoint, and Hermitian properties
Here we report a few results concerning linearity, adjoint and Hermitian character in the quaternionic setting.
Lemma 4.2 ($\mathbb{C}_{\mathbf{i}}$-linearity of sandwich Fourier coefficients).
Fix an ordered partition $S_{L}\dot{\cup}S_{R}=[d]$. For $F,G\in L^{1}(\mathbb{T}^{d};\mathbb{H}^{s\times t})$ and $a,b\in\mathbb{C}_{\mathbf{i}}$, define
$\widehat{F}^{(S_{L},S_{R})}(\bm{k}):=(2\pi)^{-d}\!\int_{\mathbb{T}^{d}}e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle_{S_{L}}}\,F(\bm{\theta})\,e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle_{S_{R}}}\,d\bm{\theta},\qquad\bm{k}\in\mathbb{Z}^{d}.$ (4.8)
Then, for all $\bm{k}\in\mathbb{Z}^{d}$,
$\widehat{\,aF+bG\,}^{(S_{L},S_{R})}(\bm{k})=a\,\widehat{F}^{(S_{L},S_{R})}(\bm{k})+b\,\widehat{G}^{(S_{L},S_{R})}(\bm{k}).$ (4.9)
Proof.
Since the kernels $e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle_{S_{L}}}$ and $e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle_{S_{R}}}$ take values in $\mathbb{C}_{\mathbf{i}}$, they commute with $a,b$. Using only associativity in $\mathbb{H}$, commutation with $\mathbb{C}_{\mathbf{i}}$-valued kernels, and the $\mathbb{R}$-linearity of the Lebesgue integral, we get
$\widehat{\,aF+bG\,}^{(S_{L},S_{R})}(\bm{k})$ $=(2\pi)^{-d}\!\int e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle_{S_{L}}}\,(aF+bG)\,e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle_{S_{R}}}\,d\bm{\theta}$
$=(2\pi)^{-d}\!\int\bigl(a\,e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle_{S_{L}}}Fe^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle_{S_{R}}}+b\,e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle_{S_{L}}}Ge^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle_{S_{R}}}\bigr)\,d\bm{\theta}$
$=a\,\widehat{F}^{(S_{L},S_{R})}(\bm{k})+b\,\widehat{G}^{(S_{L},S_{R})}(\bm{k}),$
which proves (4.9). ∎
Proposition 4.3 ($\mathbb{C}_{\mathbf{i}}$-linearity of $d$-level quaternion Toeplitz matrices).
Let $T_{\bm{n}}^{(S_{L},S_{R})}(F)\in\mathbb{H}^{(N_{\bm{n}}s)\times(N_{\bm{n}}t)}$ be defined by
$\bigl[T_{\bm{n}}^{(S_{L},S_{R})}(F)\bigr]_{\alpha,\beta}=\widehat{F}^{(S_{L},S_{R})}(\bm{k}),\qquad\bm{k}:=\alpha-\beta,\ \ \alpha,\beta\in\Lambda_{\bm{n}},$
with $\Lambda_{\bm{n}}$ as in (4.1). Then, for all $a,b\in\mathbb{C}_{\mathbf{i}}$ and $F,G\in L^{1}(\mathbb{T}^{d};\mathbb{H}^{s\times t})$,
$T_{\bm{n}}^{(S_{L},S_{R})}(aF+bG)=a\,T_{\bm{n}}^{(S_{L},S_{R})}(F)+b\,T_{\bm{n}}^{(S_{L},S_{R})}(G).$ (4.10)
Proof.
By definition, each block entry of $T_{\bm{n}}^{(S_{L},S_{R})}(H)$ equals $\widehat{H}^{(S_{L},S_{R})}(\bm{k})$ with $\bm{k}=\alpha-\beta$. Applying Lemma 4.2 entrywise with $H=aF+bG$ gives (4.10). ∎
Now, we characterize the adjoints of quaternion Toeplitz matrices and we provide criteria for checking the Hermitian character. Consider
$F:\mathbb{T}^{d}\to\mathbb{H}^{s\times t},\qquad F(\bm{\theta})=Z(\bm{\theta})+W(\bm{\theta})\mathbf{j},$
with $Z,W\in L^{1}(\mathbb{T}^{d};\mathbb{C}_{\mathbf{i}}^{\,s\times t})$. Note that for any ordered partition $S_{L}\dot{\cup}S_{R}=[d]$ and $\bm{k}=(\bm{k}_{L},\bm{k}_{R})$ supported on $S_{L},S_{R}$ we have
$\widehat{F}^{(S_{L},S_{R})}(\bm{k})=\widehat{Z}(\,\bm{k}_{L}+\bm{k}_{R}\,)+\widehat{W}(\,\bm{k}_{L}-\bm{k}_{R}\,)\,\mathbf{j},$ (4.11)
where the hats denote classical complex Fourier coefficients.
Theorem 4.4 (Adjoint identities).
Let $F:\mathbb{T}^{d}\to\mathbb{H}^{s\times t}$ with $F(\bm{\theta})=Z(\bm{\theta})+W(\bm{\theta})\mathbf{j}$, $Z,W\in L^{1}(\mathbb{T}^{d};\mathbb{C}_{\mathbf{i}}^{\,s\times t})$, and let $\bm{n}\in\mathbb{N}^{d}$. Then
$\bigl(T_{\bm{n}}^{(L)}(F)\bigr)^{*}=T_{\bm{n}}^{(R)}(F^{*}),\qquad\bigl(T_{\bm{n}}^{(R)}(F)\bigr)^{*}=T_{\bm{n}}^{(L)}(F^{*}),\qquad\bigl(T_{\bm{n}}^{(S_{L},S_{R})}(F)\bigr)^{*}=T_{\bm{n}}^{(S_{R},S_{L})}(F^{*}).$ (4.12)
Proof.
By the *-closure of integrals,
$\bigl(\widehat{F}^{(L)}(-\bm{k})\bigr)^{*}=\frac{1}{(2\pi)^{d}}\!\int_{\mathbb{T}^{d}}\bigl(F(\bm{\theta})\bigr)^{*}\,e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle}\,d\bm{\theta}=\widehat{(F^{*})}^{(R)}(\bm{k}),$
and similarly $\bigl(\widehat{F}^{(R)}(-\bm{k})\bigr)^{*}=\widehat{(F^{*})}^{(L)}(\bm{k})$. For the two-sided case, writing $\bm{k}=(\bm{k}_{L},\bm{k}_{R})$ supported on $S_{L},S_{R}$,
$\bigl(\widehat{F}^{(S_{L},S_{R})}(-\bm{k}_{L},-\bm{k}_{R})\bigr)^{*}=\widehat{(F^{*})}^{(S_{R},S_{L})}(\bm{k}_{R},\bm{k}_{L}).$
Hence, for $\alpha,\beta\in\Lambda_{\bm{n}}$,
$\bigl(T_{\bm{n}}^{(\tau)}(F)\bigr)^{*}_{\alpha\beta}=\bigl(\widehat{F}^{(\tau)}(\beta-\alpha)\bigr)^{*}=\widehat{(F^{*})}^{(\tau^{\prime})}(\alpha-\beta)=\bigl(T_{\bm{n}}^{(\tau^{\prime})}(F^{*})\bigr)_{\alpha\beta},$
where $(\tau,\tau^{\prime})$ is either $(L,R)$, $(R,L)$, or $(S_{L},S_{R})$, $(S_{R},S_{L})$. This proves the claim. ∎
We now generalize the Hermitian character proved in [35] for unilevel scalar right Toeplitz matrices. As in the basic case, the Hermitian character is completely determined by the Fourier coefficient behavior of the generating functions. In particular, the following result holds true.
Theorem 4.5 (Hermitian character for single-axis quaternion Toeplitz matrices).
Assume $s=t$ and let $F(\bm{\theta})=Z(\bm{\theta})+W(\bm{\theta})\mathbf{j}$ with $Z,W\in L^{1}(\mathbb{T}^{d};\mathbb{C}_{\mathbf{i}}^{\,s\times s})$. Fix an ordered partition $S_{L}\dot{\cup}S_{R}=[d]$. For $\bm{n}\in\mathbb{N}^{d}$ set
$T_{\bm{n}}^{(S_{L},S_{R})}(F)=\bigl[\widehat{F}^{(S_{L},S_{R})}(\alpha-\beta)\bigr]_{\alpha,\beta\in\Lambda_{\bm{n}}}.$
Then the following are equivalent:
  1. (i)
    $T_{\bm{n}}^{(S_{L},S_{R})}(F)$ is Hermitian for all $\bm{n}$;
  2. (ii)
    $Z(\bm{\theta})$ is essentially Hermitian on $\mathbb{T}^{d}$ and $W(-\bm{\theta})=-\,W(\bm{\theta})^{\mathsf{T}}$ a.e. on $\mathbb{T}^{d}$.
In particular, the claim does not depend on the choice of the ordered partition.
Proof.
For a fixed $\bm{n}$, stating that $T_{\bm{n}}^{(S_{L},S_{R})}(F)$ is Hermitian is equivalent to
$\widehat{F}^{(S_{L},S_{R})}(\bm{k})=\bigl(\widehat{F}^{(S_{L},S_{R})}(-\bm{k})\bigr)^{*}\qquad\text{for all }\bm{k}\in\Delta(\bm{n}),$ (4.13)
where $\Delta(\bm{n})=\{\alpha-\beta:\ \alpha,\beta\in\Lambda_{\bm{n}}\}$. Since the statement holds for all $\bm{n}$, (4.13) is equivalent to the same identity for all $\bm{k}\in\mathbb{Z}^{d}$.
Write $\bm{k}=(\bm{k}_{L},\bm{k}_{R})$ supported on $S_{L},S_{R}$. By (4.11),
$\widehat{F}^{(S_{L},S_{R})}(\bm{k})=\widehat{Z}(\bm{k}_{L}+\bm{k}_{R})+\widehat{W}(\bm{k}_{L}-\bm{k}_{R})\,\mathbf{j}.$
Set
$\bm{p}:=\bm{k}_{L}+\bm{k}_{R},\qquad\bm{q}:=\bm{k}_{L}-\bm{k}_{R}\quad(\in\mathbb{Z}^{d}).$
Then (4.13) becomes, for all $\bm{p},\bm{q}\in\mathbb{Z}^{d}$,
$\widehat{Z}(\bm{p})+\widehat{W}(\bm{q})\,\mathbf{j}=\Bigl(\widehat{Z}(-\bm{p})+\widehat{W}(-\bm{q})\,\mathbf{j}\Bigr)^{*}=\widehat{Z}(-\bm{p})^{*}+(\widehat{W}(-\bm{q})\,\mathbf{j})^{*}.$
Using $\mathbf{j}^{*}=-\mathbf{j}$ and $\mathbf{j}A=\overline{A}\,\mathbf{j}$ for $A\in\mathbb{C}_{\mathbf{i}}^{\,s\times s}$, together with $\overline{A^{*}}=A^{\mathsf{T}}$, we obtain
$(\widehat{W}(-\bm{q})\,\mathbf{j})^{*}=\mathbf{j}^{*}\widehat{W}(-\bm{q})^{*}=-\mathbf{j}\,\widehat{W}(-\bm{q})^{*}=-\,\overline{\widehat{W}(-\bm{q})^{*}}\,\mathbf{j}=-\,\bigl(\widehat{W}(-\bm{q})\bigr)^{\mathsf{T}}\,\mathbf{j}.$
Hence
$\widehat{Z}(\bm{p})+\widehat{W}(\bm{q})\,\mathbf{j}=\widehat{Z}(-\bm{p})^{*}-\bigl(\widehat{W}(-\bm{q})\bigr)^{\mathsf{T}}\,\mathbf{j}.$
Since $1$ and $\mathbf{j}$ are linearly independent over the slice $\mathbb{C}_{\mathbf{i}}$, this is equivalent to
$\widehat{Z}(\bm{p})=\widehat{Z}(-\bm{p})^{*}\qquad\text{and}\qquad\widehat{W}(\bm{q})=-\,\bigl(\widehat{W}(-\bm{q})\bigr)^{\mathsf{T}}\quad\text{for all }\bm{p},\bm{q}\in\mathbb{Z}^{d}.$ (4.14)
[…]
Finally, the derivation uses only (4.11) and the fact that $\bm{p}$ and $\bm{q}$ range over all of $\mathbb{Z}^{d}$ as $\bm{k}$ varies, so the conclusion is independent of the ordered partition. ∎
4.3 Symplectic embedding of Toeplitz matrices and spectral/singular value distributions of Toeplitz matrix sequences
The aim of the present section is to derive spectral/singular value distributions for block multilevel Toeplitz matrix sequences in the quaternion setting. We start by defining the proper embeddings.
4.3.1 Left/right and sandwich embeddings
Throughout this section, let $F=Z+W\mathbf{j}\in L^{1}(\mathbb{T}^{d};\mathbb{H}^{s\times t})$ with $Z,W:\mathbb{T}^{d}\to\mathbb{C}_{\mathbf{i}}^{\,s\times t}$ slice-valued. Let $\Phi:\mathbb{H}^{s\times t}\to\mathbb{C}^{(2s)\times(2t)}$ be the block symplectic embedding on $s\times t$ matrices
$\Phi(Z+W\mathbf{j})=\begin{bmatrix}Z&W\\[2.0pt] -\overline{W}&\overline{Z}\end{bmatrix},\qquad Z,W\in\mathbb{C}_{\mathbf{i}}^{\,s\times t}.$
If $A\in\mathbb{H}^{N_{\bm{n}}s\times N_{\bm{n}}t}$ is an $s\times t$-block matrix, we apply $\Phi$ $s\times t$-blockwise, meaning
$\Phi(A):=\bigl[\Phi(A_{\alpha\beta})\bigr]_{\alpha,\beta\in{\Lambda_{\bm{n}}}}\in\mathbb{C}^{(2sN_{\bm{n}})\times(2tN_{\bm{n}})},$
with $A_{\alpha,\beta}\in\mathbb{H}^{s\times t}$. We will use the following identity
$\Phi(\lambda A\rho)=\operatorname{diag}(\lambda,\overline{\lambda})\,\Phi(A)\,\operatorname{diag}(\rho,\overline{\rho})\quad\bigl(\lambda\in\mathbb{C}_{\mathbf{i}}^{\,s\times s},\ \rho\in\mathbb{C}_{\mathbf{i}}^{\,t\times t}\bigr),$ (4.15)
where $\operatorname{diag}(\lambda,\overline{\lambda})$ and $\operatorname{diag}(\rho,\overline{\rho})$ are $2\times 2$ block-diagonal matrices with blocks of sizes $s\times s$ and $t\times t$, respectively. When $\lambda,\rho\in\mathbb{C}_{\mathbf{i}}$ are scalars we read them as $\lambda I_{s}$ and $\rho I_{t}$.
For a complex matrix symbol $G:\mathbb{T}^{d}\to\mathbb{C}^{(2s)\times(2t)}$, we write
$\widehat{G}(\bm{k})=(2\pi)^{-d}\!\int_{\mathbb{T}^{d}}G(\bm{\theta})\,e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle}\,d\bm{\theta},\qquad T_{\bm{n}}(G):=\bigl[\widehat{G}(\alpha-\beta)\bigr]_{\alpha,\beta\in\Lambda_{\bm{n}}}.$
Theorem 4.6 (Left/right embedding).
With the decomposition $F(\bm{\theta})=Z(\bm{\theta})+W(\bm{\theta})\mathbf{j}$ from Lemma 2.15, define the complex symbols
$G_{\mathrm{L}}(\bm{\theta})=\begin{pmatrix}Z(\bm{\theta})&W(\bm{\theta})\\[2.0pt] -\overline{W(-\bm{\theta})}&\overline{Z(-\bm{\theta})}\end{pmatrix},\qquad G_{\mathrm{R}}(\bm{\theta})=\begin{pmatrix}Z(\bm{\theta})&W(-\bm{\theta})\\[2.0pt] -\overline{W(\bm{\theta})}&\overline{Z(-\bm{\theta})}\end{pmatrix}$.
Then, for every $\bm{n}\in\mathbb{N}^{d}$,
$\Phi\!\bigl(T_{\bm{n}}^{(\mathrm{L})}(F)\bigr)=T_{\bm{n}}(G_{\mathrm{L}}),\qquad\Phi\!\bigl(T_{\bm{n}}^{(\mathrm{R})}(F)\bigr)=T_{\bm{n}}(G_{\mathrm{R}}).$
Proof.
Let $\bm{k}=\alpha-\beta\in\mathbb{Z}^{d}$. Using (4.15) with the left multiplier $\lambda(\bm{\theta})=e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle}I_{s}$ for the left case (respectively the right multiplier $\rho(\bm{\theta})=e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle}I_{t}$ for the right case), we obtain
$\Phi\!\bigl(\widehat{F}^{(\mathrm{L})}(\bm{k})\bigr)=\frac{1}{(2\pi)^{d}}\!\int_{\mathbb{T}^{d}}\!\operatorname{diag}\!\bigl(e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle}I_{s},\,e^{+\mathbf{i}\langle\bm{k},\bm{\theta}\rangle}I_{s}\bigr)\,\Phi\!\bigl(F(\bm{\theta})\bigr)\,d\bm{\theta},$
and
$\Phi\!\bigl(\widehat{F}^{(\mathrm{R})}(\bm{k})\bigr)=\frac{1}{(2\pi)^{d}}\!\int_{\mathbb{T}^{d}}\!\Phi\!\bigl(F(\bm{\theta})\bigr)\,\operatorname{diag}\!\bigl(e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle}I_{t},\,e^{+\mathbf{i}\langle\bm{k},\bm{\theta}\rangle}I_{t}\bigr)\,d\bm{\theta}.$
Since $\Phi(F)=\bigl[\begin{smallmatrix}Z&W\\ -\overline{W}&\overline{Z}\end{smallmatrix}\bigr]$, a direct multiplication of the integrands yields the phase factors $Z\,e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle}$ and $W\,e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle}$ in the left case, and $Z\,e^{-\mathbf{i}\langle\bm{k},\bm{\theta}\rangle}$ and $W\,e^{+\mathbf{i}\langle\bm{k},\bm{\theta}\rangle}$ in the right case; the lower row carries the corresponding conjugate phases. Therefore
$\Phi\!\bigl(\widehat{F}^{(\mathrm{L})}(\bm{k})\bigr)=\begin{pmatrix}\widehat{Z}(\bm{k})&\widehat{W}(\bm{k})\\[2.0pt] -\overline{\widehat{W}(\bm{k})}&\overline{\widehat{Z}(\bm{k})}\end{pmatrix},\qquad\Phi\!\bigl(\widehat{F}^{(\mathrm{R})}(\bm{k})\bigr)=\begin{pmatrix}\widehat{Z}(\bm{k})&\widehat{W}(-\bm{k})\\[2.0pt] -\overline{\widehat{W}(-\bm{k})}&\overline{\widehat{Z}(\bm{k})}\end{pmatrix}$.
These are precisely the $\bm{k}$-th Fourier coefficients of $G_{\mathrm{L}}$ and $G_{\mathrm{R}}$, respectively. Setting $\bm{k}=\alpha-\beta$ gives the Toeplitz identities. ∎
We now treat the two-sided sandwich case using the involution $r_{S_{L}}:\mathbb{T}^{d}\to\mathbb{T}^{d}$ from Proposition 4.1. Note that $r_{S_{L}}\circ(-\mathrm{id})$ flips the $S_{R}$-coordinates and that $r_{S_{L}}(-\bm{\theta})=r_{S_{L}}(\bm{\theta})$.
Theorem 4.7 (Sandwich embedding).
Let $F(\bm{\theta})=Z(\bm{\theta})+W(\bm{\theta})\mathbf{j}$ with $Z,W:\mathbb{T}^{d}\to\mathbb{C}_{\mathbf{i}}^{\,s\times t}$ and fix an ordered partition $S_{L}\dot{\cup}S_{R}=[d]$. Define the complex symbol
$G_{S_{L},S_{R}}(\bm{\theta})=\begin{pmatrix}Z(\bm{\theta})&W\!\bigl(r_{S_{L}}(-\bm{\theta})\bigr)\\[2.0pt] -\overline{W\!\bigl(r_{S_{L}}(\bm{\theta})\bigr)}&\overline{Z(-\bm{\theta})}\end{pmatrix}$.
Then, for every $\bm{n}\in\mathbb{N}^{d}$,
$\Phi\!\bigl(T_{\bm{n}}^{(S_{L},S_{R})}(F)\bigr)=T_{\bm{n}}\!\bigl(G_{S_{L},S_{R}}\bigr).$ (4.16)
In particular, for $S_{L}=[d],S_{R}=\varnothing$ one recovers
$G_{\mathrm{L}}(\bm{\theta})=\begin{pmatrix}Z(\bm{\theta})&W(\bm{\theta})\\[2.0pt] -\overline{W(-\bm{\theta})}&\overline{Z(-\bm{\theta})}\end{pmatrix},\qquad\text{and for }S_{L}=\varnothing,S_{R}=[d]\text{ one gets }G_{\mathrm{R}}(\bm{\theta})=\begin{pmatrix}Z(\bm{\theta})&W(-\bm{\theta})\\[2.0pt] -\overline{W(\bm{\theta})}&\overline{Z(-\bm{\theta})}\end{pmatrix}$.
Proof.
Let $\bm{k}=\alpha-\beta\in\mathbb{Z}^{d}$ and write $\bm{k}=(\bm{k}_{L},\bm{k}_{R})$ supported on $S_{L},S_{R}$. By the definition of the sandwich Fourier coefficients,
$\widehat{F}^{(S_{L},S_{R})}(\bm{k})=\frac{1}{(2\pi)^{d}}\!\int_{\mathbb{T}^{d}}e^{-\mathbf{i}\langle\bm{k}_{L},\bm{\theta}\rangle}\,(Z+W\mathbf{j})\,e^{-\mathbf{i}\langle\bm{k}_{R},\bm{\theta}\rangle}\,d\bm{\theta}.$
Using the covariance identity for the symplectic embedding (valid for $\lambda,\rho\in\mathbb{C}_{\mathbf{i}}$)
$\Phi(\lambda A\rho)=\operatorname{diag}(\lambda,\overline{\lambda})\,\Phi(A)\,\operatorname{diag}(\rho,\overline{\rho}),$
with
$\lambda(\bm{\theta})=e^{-\mathbf{i}\langle\bm{k}_{L},\bm{\theta}\rangle}I_{s},\qquad\rho(\bm{\theta})=e^{-\mathbf{i}\langle\bm{k}_{R},\bm{\theta}\rangle}I_{t},$
we obtain
$\Phi\!\bigl(\widehat{F}^{(S_{L},S_{R})}(\bm{k})\bigr)=\frac{1}{(2\pi)^{d}}\!\int_{\mathbb{T}^{d}}\begin{bmatrix}e^{-\mathbf{i}\langle\bm{k}_{L},\bm{\theta}\rangle}I_{s}&0\\[1.0pt] 0&e^{+\mathbf{i}\langle\bm{k}_{L},\bm{\theta}\rangle}I_{s}\end{bmatrix}\begin{bmatrix}Z&W\\[1.0pt] -\overline{W}&\overline{Z}\end{bmatrix}\begin{bmatrix}e^{-\mathbf{i}\langle\bm{k}_{R},\bm{\theta}\rangle}I_{t}&0\\[1.0pt] 0&e^{+\mathbf{i}\langle\bm{k}_{R},\bm{\theta}\rangle}I_{t}\end{bmatrix}\,d\bm{\theta}.$
Multiplying the integrand yields the four entries
$Z\,e^{-\mathbf{i}\langle\bm{k}_{L}+\bm{k}_{R},\bm{\theta}\rangle},\quad W\,e^{-\mathbf{i}\langle\bm{k}_{L}-\bm{k}_{R},\bm{\theta}\rangle},\quad-\overline{W}\,e^{+\mathbf{i}\langle\bm{k}_{L}-\bm{k}_{R},\bm{\theta}\rangle},\quad\overline{Z}\,e^{+\mathbf{i}\langle\bm{k}_{L}+\bm{k}_{R},\bm{\theta}\rangle}.$
Hence
$\Phi\!\bigl(\widehat{F}^{(S_{L},S_{R})}(\bm{k})\bigr)=\begin{pmatrix}\widehat{Z}(\bm{k}_{L}+\bm{k}_{R})&\widehat{W}(\bm{k}_{L}-\bm{k}_{R})\\[2.0pt] -\overline{\widehat{W}(\bm{k}_{L}-\bm{k}_{R})}&\overline{\widehat{Z}(\bm{k}_{L}+\bm{k}_{R})}\end{pmatrix}$.
By standard changes of variables on $\mathbb{T}^{d}$,
$\widehat{H\circ(-\mathrm{id})}(\bm{k})=\widehat{H}(-\bm{k}),\qquad\widehat{H\circ r_{S_{L}}}(\bm{k})=\widehat{H}\!\bigl(r_{S_{L}}(\bm{k})\bigr),$
and since $r_{S_{L}}(-\bm{k})=(\bm{k}_{L},-\bm{k}_{R})$, we have
$\widehat{W\circ r_{S_{L}}\circ(-\mathrm{id})}(\bm{k})=\widehat{W}\!\bigl(r_{S_{L}}(-\bm{k})\bigr)=\widehat{W}(\bm{k}_{L}-\bm{k}_{R}),\qquad\widehat{\overline{Z(-\cdot)}}(\bm{k})=\overline{\widehat{Z}(\bm{k})}.$
Therefore the last displayed matrix equals $\widehat{G_{S_{L},S_{R}}}(\bm{k})$. Setting $\bm{k}=\alpha-\beta$ and assembling by blocks yields $\Phi\!\bigl(T_{\bm{n}}^{(S_{L},S_{R})}(F)\bigr)=T_{\bm{n}}\!\bigl(G_{S_{L},S_{R}}\bigr),$ which is (4.16). ∎
4.3.2 Singular value and spectral distribution via the embedding approach
In this part, we use the embedding identities proved in the previous sections to transfer spectral and singular value distribution results known for complex block Toeplitz sequences to quaternion Toeplitz sequences. Throughout, write $\bm{n}=(n_{1},\dots,n_{d})\in\mathbb{N}^{d}$ and $N_{\bm{n}}=\prod_{\ell=1}^{d}n_{\ell}$. For $\tau\in\{L,R,(S_{L},S_{R})\}$ we associate the complex symbol
$G_{\tau}(\bm{\theta}):=\begin{cases}G_{\mathrm{L}}(\bm{\theta}),&\tau=L,\\[2.0pt] G_{\mathrm{R}}(\bm{\theta}),&\tau=R,\\[2.0pt] G_{S_{L},S_{R}}(\bm{\theta}),&\tau=(S_{L},S_{R}),\end{cases}$
where $G_{\mathrm{L}},G_{\mathrm{R}},G_{S_{L},S_{R}}$ are those in Theorems 4.6 and 4.7.
Theorem 4.8 (Singular value distribution and Hermitian spectral distribution).
Let $F(\bm{\theta})=Z(\bm{\theta})+W(\bm{\theta})\mathbf{j}$ with $Z,W\in L^{1}(\mathbb{T}^{d};\mathbb{C}_{\mathbf{i}}^{\,s\times t})$. For each $\tau\in\{L,R,(S_{L},S_{R})\}$,
$\bigl\{\,T_{\bm{n}}^{(\tau)}(F)\,\bigr\}_{\bm{n}}\ \sim_{\sigma}\ G_{\tau}.$ (4.16)
If, in addition, $s=t$ and $T_{\bm{n}}^{(\tau)}(F)$ is Hermitian for all $\bm{n}$, then also
$\bigl\{\,T_{\bm{n}}^{(\tau)}(F)\,\bigr\}_{\bm{n}}\ \sim_{\lambda}\ G_{\tau}.$
Proof.
By the embedding identities established above,
$\Phi\!\bigl(T_{\bm{n}}^{(L)}(F)\bigr)=T_{\bm{n}}(G_{\mathrm{L}}),\qquad\Phi\!\bigl(T_{\bm{n}}^{(R)}(F)\bigr)=T_{\bm{n}}(G_{\mathrm{R}}),\qquad\Phi\!\bigl(T_{\bm{n}}^{(S_{L},S_{R})}(F)\bigr)=T_{\bm{n}}(G_{S_{L},S_{R}}).$
Hence it suffices to analyze $T_{\bm{n}}(G_{\tau})$ on the complex side. The singular value distribution for complex block Toeplitz matrix sequences with $L^{1}$ matrix symbols is an established result (see, e.g., [20, 21, 5]), thus we have $\{T_{\bm{n}}(G_{\tau})\}_{\bm{n}}\sim_{\sigma}G_{\tau}$. By Corollary 2.9, the singular values of $\Phi(A)$ are those of $A$, each duplicated, so the quaternion-side normalization matches the complex one, giving (4.16).
If $s=t$ and every $T_{\bm{n}}^{(\tau)}(F)$ is Hermitian, then $\Phi\!\bigl(T_{\bm{n}}^{(\tau)}(F)\bigr)$ is Hermitian as well. The complex eigenvalue distribution for Hermitian Toeplitz sequences with $L^{1}$ symbols (again [20, 21, 5]) gives $\{T_{\bm{n}}(G_{\tau})\}_{\bm{n}}\sim_{\lambda}G_{\tau}$. The embedding argument transfers the claim back to the quaternion sequence. ∎
For the next result, we need to pose hypotheses on the topology of spectral essential range of the symbols, in accordance with definition below.
Definition 4.9.
Let $H:\mathbb{T}^{d}\to\mathbb{C}^{m\times m}$ be a measurable matrix-valued function. For each $j$, let $R(\lambda_{j}(H))$ denote the essential range of the scalar function $\bm{\theta}\mapsto\lambda_{j}(H(\bm{\theta}))$ (with any measurable labeling of eigenvalues). Then the spectral essential range of $H$ is
$R(H):=\bigcup_{j=1}^{m}R\bigl(\lambda_{j}(H)\bigr).$
This set is clearly independent of the labeling.
Under mild topological assumptions on the spectral essential range of the embedding symbols, known as Tilli conditions [61, 13], we can go beyond the Hermitian case and describe the (canonical) eigenvalue distribution of non-Hermitian Toeplitz sequences generated by bounded functions.
Theorem 4.10 (Spectral distribution for Tilli-class symbols).
Assume $s=t$ and $F(\bm{\theta})=Z(\bm{\theta})+W(\bm{\theta})\mathbf{j}$ with $Z,W\in L^{\infty}(\mathbb{T}^{d};\mathbb{C}_{\mathbf{i}}^{\,s\times s})$. For $\tau\in\{L,R,(S_{L},S_{R})\}$, consider the spectral essential range $R(G_{\tau})$. Suppose that $G_{\tau}$ belongs to the Tilli class, i.e.,
$\qquad\mathrm{int}\,R(G_{\tau})=\varnothing\ \ \text{and}\ \ \mathbb{C}\setminus R(G_{\tau})\ \text{is connected}.$
Then
$\bigl\{\,T_{\bm{n}}^{(\tau)}(F)\,\bigr\}_{\bm{n}}\ \sim_{\lambda}\ G_{\tau}.$ (4.17)
Proof.
By the embedding identities, $\Phi\!\bigl(T_{\bm{n}}^{(\tau)}(F)\bigr)=T_{\bm{n}}(G_{\tau})$. Since $G_{\tau}\in L^{\infty}(\mathbb{T}^{d};\mathbb{C}^{(2s)\times(2s)})$ and it belongs to the complex Tilli class, the complex Tilli theorem (see [13] for the block multilevel version used here) yields $\{T_{\bm{n}}(G_{\tau})\}_{\bm{n}}\sim_{\lambda}G_{\tau}$. The eigenvalue distribution for the quaternion sequence follows from the embedding. ∎
Remark 4.11 (The canonical essential range satifies the Tilli class).
Note that
4.4 Schatten–norm estimates
In this section we derive Schatten bounds for quaternionic multilevel Toeplitz matrices via the complex symplectic embedding. Let $F(\bm{\theta})=Z(\bm{\theta})+W(\bm{\theta})\mathbf{j}$ with $Z,W\in L^{p}(\mathbb{T}^{d};\mathbb{C}_{\mathbf{i}}^{\,s\times t})$ and $1\leq p\leq\infty$. We use the embedded complex symbols from Theorems 4.6 and 4.7, denoted $G_{\tau}(\bm{\theta})$ with $\tau\in\{L,R,(S_{L},S_{R})\}$. For the next results we recall the Schatten properties
$\|\Phi(X)\|_{p}=2^{1/p}\,\|X\|_{p}\quad\text{(with the convention $2^{1/\infty}=1$),}$
and,
$\|\Phi\!\circ F\|_{L^{p}(\mathbb{T}^{d};\mathbb{C}^{(2s)\times(2t)})}=2^{1/p}\,\|F\|_{L^{p}(\mathbb{T}^{d};\mathbb{H}^{s\times t})}.$ (4.18)
Lemma 4.12 (Complex matrix-valued Toeplitz Schatten bound).
Let $G\in L^{p}(\mathbb{T}^{d};\mathbb{C}^{a\times b})$ with $1\leq p\leq\infty$ and $\bm{n}\in\mathbb{N}^{d}$, $N_{\bm{n}}:=\prod_{\ell=1}^{d}n_{\ell}$. Then
$\bigl\|T_{\bm{n}}(G)\bigr\|_{p}\ \leq\ \left(\frac{N_{\bm{n}}}{(2\pi)^{d}}\right)^{\!1/p}\ \|G\|_{L^{p}(\mathbb{T}^{d};\mathbb{C}^{a\times b})}.$ (4.19)
Moreover, the constant $1$ is optimal.
Theorem 4.13 (Quaternionic matrix-valued Toeplitz–Schatten bound).
Let $1\leq p\leq\infty$ and $F\in L^{p}(\mathbb{T}^{d};\mathbb{H}^{s\times t})$. For any $\tau\in\{L,R,(S_{L},S_{R})\}$ and $\bm{n}\in\mathbb{N}^{d}$, with $N_{\bm{n}}:=\prod_{\ell=1}^{d}n_{\ell}$,
$\bigl\|T^{(\tau)}_{\bm{n}}(F)\bigr\|_{p}\ \leq\ {4}\left(\frac{N_{\bm{n}}}{(2\pi)^{d}}\right)^{\!1/p}\,\|F\|_{L^{p}(\mathbb{T}^{d};\mathbb{H}^{s\times t})}.$ (4.20)
Proof.
We consider a general sandwich-kernel setting, thus consider the embedding symbol
$G_{S_{L},S_{R}}(\bm{\theta})=\begin{pmatrix}Z(\bm{\theta})&W\!\bigl(r_{S_{L}}(-\bm{\theta})\bigr)\\[2.0pt] -\overline{W\!\bigl(r_{S_{L}}(\bm{\theta})\bigr)}&\overline{Z(-\bm{\theta})}\end{pmatrix}$,
where $r_{S_{L}}:\mathbb{T}^{d}\to\mathbb{T}^{d}$ is the reflection of left-kernel variables. Define the row projections
$P_{1}:=\operatorname{diag}(I_{s},0)\in\mathbb{C}^{(2s)\times(2s)},\qquad P_{2}:=\operatorname{diag}(0,I_{s}),$
and the respective column projections
$E_{1}:=\operatorname{diag}(I_{t},0)\in\mathbb{C}^{(2t)\times(2t)},\qquad E_{2}:=\operatorname{diag}(0,I_{t}).$
Clearly, $\|P_{j}\|_{\infty}=\|E_{j}\|_{\infty}=1$ for $j=1,2$.
Using the projections, we decompose $G_{S_{L},S_{R}}(\bm{\theta})$ into the sum of four terms:
$G_{S_{L},S_{R}}(\bm{\theta})$ $=P_{1}\,\Phi\!\bigl(F(\bm{\theta})\bigr)\,E_{1}+P_{1}\,\Phi\!\bigl(F(r_{S_{L}}(-\bm{\theta}))\bigr)\,E_{2}$
$\quad+P_{2}\,\Phi\!\bigl(F(r_{S_{L}}(\bm{\theta}))\bigr)\,E_{1}+P_{2}\,\Phi\!\bigl(F(-\bm{\theta})\bigr)\,E_{2},$
where, explicitily,
$P_{1}\,\Phi\!\bigl(F(\bm{\theta})\bigr)E_{1}=\begin{pmatrix}Z(\bm{\theta})&0\\[2.0pt] 0&0\end{pmatrix},\quad P_{1}\,\Phi\!\bigl(F(r_{S_{L}}(-\bm{\theta}))\bigr)E_{2}=\begin{pmatrix}0&W(r_{S_{L}}(-\bm{\theta}))\\[2.0pt] 0&0\end{pmatrix}$,
$P_{2}\,\Phi\!\bigl(F(r_{S_{L}}(\bm{\theta}))\bigr)E_{1}=\begin{pmatrix}0&0\\[2.0pt] -\overline{W(r_{S_{L}}(\bm{\theta}))}&0\end{pmatrix},\quad P_{2}\,\Phi\!\bigl(F(-\bm{\theta})\bigr)E_{2}=\begin{pmatrix}0&0\\[2.0pt] 0&\overline{Z(-\bm{\theta})}\end{pmatrix}$.
Now for any choice of possible contractions, we have
$\|AXB\|_{p}\leq\|A\|\,\|X\|_{p}\,\|B\|\leq\|X\|_{p}.$
Hence, by using (4.4), for any $\bm{\theta}$ we deduce
$\|G_{S_{L},S_{R}}(\bm{\theta})\|_{p}$ $\leq\sum_{j=1}^{4}\bigl\|\Phi\!\bigl(F(\varphi_{j}(\bm{\theta}))\bigr)\bigr\|_{p},$
where $\varphi_{j}(\bm{\theta})$ is either $\bm{\theta}\mapsto\bm{\theta}$, $\bm{\theta}\mapsto-\bm{\theta}$, or $\bm{\theta}\mapsto r_{S_{L}}(\bm{\theta})$. Note that all of these functions are measure preserving and are involution in $\mathbb{T}^{d}$. Hence, passing to the $L^{p}$ norm, we have
$\bigl\|\Phi\circ F\circ\varphi_{j}\bigr\|_{L^{p}}=\|\Phi\circ F\|_{L^{p}}.$
Summing all the terms we obtain
$\|G_{\tau}\|_{L^{p}(\mathbb{T}^{d})}\leq 4\,\|\Phi\circ F\|_{L^{p}(\mathbb{T}^{d})}=4\cdot 2^{1/p}\,\|F\|_{L^{p}(\mathbb{T}^{d};\mathbb{H}^{s\times t})},$
by the symplectic $L^{p}$ isometry.
Furthermore, we consider the corresponding Toeplitz matrix $T_{\bm{n}}^{(S_{L},S_{R})}(F)$. By assumption, we have
$\Phi\!\bigl(T_{\bm{n}}^{(S_{L},S_{R})}(F)\bigr)=T_{\bm{n}}(G_{\tau}).$
By the Schatten scaling under $\Phi$,
$\bigl\|T_{\bm{n}}^{(\tau)}(F)\bigr\|_{p}=2^{-1/p}\bigl\|T_{\bm{n}}(G_{\tau})\bigr\|_{p}.$
Further, applying the complex Toeplitz bound (Lemma 4.12) and the $L^{p}$ bound for $G_{\tau}$, we obtain
$\bigl\|T_{\bm{n}}^{(\tau)}(F)\bigr\|_{p}$ $\leq 2^{-1/p}\left(\frac{N_{\bm{n}}}{(2\pi)^{d}}\right)^{1/p}\|G_{\tau}\|_{L^{p}(\mathbb{T}^{d})}$
$\leq 2^{-1/p}\left(\frac{N_{\bm{n}}}{(2\pi)^{d}}\right)^{1/p}\cdot 4\cdot 2^{1/p}\,\|F\|_{L^{p}(\mathbb{T}^{d};\mathbb{H}^{s\times t})}$
$=4\left(\frac{N_{\bm{n}}}{(2\pi)^{d}}\right)^{1/p}\|F\|_{L^{p}(\mathbb{T}^{d};\mathbb{H}^{s\times t})},$
which is exactly the claim we want to prove. ∎
4.5 Spectral localization of Hermitian matrices via the $2s\times 2s$ complex symbols
Lin, Pan, and Ng [35] prove strong spectral localization for unilevel scalar Hermitian right-quaternion Toeplitz matrices and provide positivity criteria. We extend these results to the multilevel and block cases via the symplectic embedding and the complex Toeplitz theory.
Throughout, let $F\in L^{1}(\mathbb{T}^{d};\mathbb{H}^{s\times s})$ and fix $\tau\in\{L,R,(S_{L},S_{R})\}$. Let $G_{\tau}$ denote the corresponding $2s\times 2s$ complex symbol (one of $G_{\mathrm{L}}$, $G_{\mathrm{R}}$, or $G_{S_{L},S_{R}}$) as in Theorems 4.6 and 4.7.
Theorem 4.14 (Spectral localization of Hermitian Toeplitz matrices).
Assume that $T_{\bm{n}}^{(\tau)}(F)$ is Hermitian for all $\bm{n}\in\mathbb{N}^{d}$. Set
$m_{\tau}:=\operatorname*{ess\,inf}_{\bm{\theta}\in\mathbb{T}^{d}}\lambda_{\min}\!\bigl(G_{\tau}(\bm{\theta})\bigr),\qquad M_{\tau}:=\operatorname*{ess\,sup}_{\bm{\theta}\in\mathbb{T}^{d}}\lambda_{\max}\!\bigl(G_{\tau}(\bm{\theta})\bigr).$
Then, for every $\bm{n}\in\mathbb{N}^{d}$,
$\sigma\!\bigl(T^{(\tau)}_{\bm{n}}(F)\bigr)\ \subset\ [\,m_{\tau},\,M_{\tau}\,].$ (4.21)
Proof.
By the embedding identities, we find
$\Phi\!\bigl(T_{\bm{n}}^{(\tau)}(F)\bigr)=T_{\bm{n}}(G_{\tau}).$
Since $T_{\bm{n}}^{(\tau)}(F)$ is Hermitian, $\Phi\!\bigl(T_{\bm{n}}^{(\tau)}(F)\bigr)$ is Hermitian and hence $T_{\bm{n}}(G_{\tau})$ is Hermitian as well. The complex spectral localization for Hermitian block Toeplitz matrices with $L^{1}$ symbols (see, e.g., [49]) gives
$\sigma\!\bigl(T_{\bm{n}}(G_{\tau})\bigr)\ \subset\ [\,m_{\tau},\,M_{\tau}\,].$
For Hermitian matrices, $\Phi$ preserves eigenvalues (it duplicates them in the complex representation but with the same values); thus (4.21) follows. ∎
Remark 4.15 (Sharpness).
In fact, we generally have open localization intervals. The closure in the extreme $m_{\tau}$ occur if and only if $\lambda_{\min}\!\bigl(G_{\tau}(\bm{\theta})\bigr)$ is essentially constant. A similar remark holds for $M_{\tau}$. The reader is referred to [49, 48] and references therein for the spectral study of block multilevel Toeplitz matrix sequences in the complex setting.
Corollary 4.16 (Positive definiteness criterion).
Assume $T_{\bm{n}}^{(\tau)}(F)$ is Hermitian for all $\bm{n}$ and let $m_{\tau}$ be as above. Then:
  1. (a)
    If $m_{\tau}>0$, then $T_{\bm{n}}^{(\tau)}(F)$ is Hermitian positive definite for every $\bm{n}$, with $\lambda_{\min}\!\bigl(T_{\bm{n}}^{(\tau)}(F)\bigr)\geq m_{\tau}$.
  2. (b)
    Conversely, if there exists $c>0$ such that $T_{\bm{n}}^{(\tau)}(F)\succeq c\,I$ for all $\bm{n}$, then $m_{\tau}\geq c$.
Proof.
(a) Apply Theorem 4.14. (b) If all eigenvalues are $\geq c$ for every $\bm{n}$, the complex spectra of $T_{\bm{n}}(G_{\tau})$ (equal to $\Phi\!\bigl(T_{\bm{n}}^{(\tau)}(F)\bigr)$) are also $\geq c$, hence $\lambda_{\min}\!\bigl(G_{\tau}(\bm{\theta})\bigr)\geq c$ a.e. by the GLT localization; thus $m_{\tau}\geq c$. ∎
4.6 Toeplitz a.c.s. approximation via circulant matrix sequences
In the current section, we recover the classical circulant approximation property for Toeplitz sequences in the quaternion setting and provide an alternative derivation of their singular value distribution using a.c.s. theory. In fact, the spectral analysis of Toeplitz sequences through circulant approximation is the “textbook” approach to this problem for two reasons: this approach is generally considered elegant and highlights in a clear way the topological structure of the space we are studying; moreover, it has a direct link with the classical preconditioning theory, where the observation that circulant matrices approximate Toeplitz ones created a huge ecosystem of fast Toeplitz preconditioning strategies (see e.g. [39] and references there reported). Numerically speaking, we expect the same link also in quaternion problems (see, for example, the fast Strang preconditioner analyzed in [35]). Barbarino and Garoni make a remark in [4], where the easy and “rigid” spectral properties of circulant matrices allow for a simpler analysis of the asymptotic spectral distribution of normal sequences of the GLT class, for which Toeplitz sequences are a subclass.
Given these premises, first, we follow the lines of [41] and decompose multilevel $s\times t$ circulant matrices under the $\mathbf{i}$-axis quaternion discrete Fourier Transform, and we extract the sparsity $X$-pattern composed of blocks of size $s\times t$ or $2s\times 2t$. Up to block permutations, this pattern provides a block-diagonalization. Then, we describe the spectrum of multilevel block quaternion circulant matrices, we show that every Toeplitz matrix of any family admits a proper circulant approximating class of sequences, and we provide an alternative approach to find the singular value or spectral distribution of Toeplitz sequences.
We recall the basic notation and tools for the Pan-Ng factorization. For $n\in\mathbb{N}$, let $A_{n}\in\{0,1\}^{n\times n}$ be the permutation matrix that maps the canonical basis in $\mathbb{H}^{n}$ according to the law $e_{s}\mapsto e_{(-s)\bmod n}$ and set $A_{\bm{n}}:=A_{n_{1}}\otimes\cdots\otimes A_{n_{d}}$. The QDFT on the one-dimensional $\mathbf{i}$-axis is defined as in the complex case (see [16]). More precisely we have
$F_{\mathbf{i},n}]_{uv}:=n^{-1/2}\exp(-2\pi\mathbf{i}\,uv/n),\qquad u,v=0,\dots,n-1,$
with the multilevel transform being $F_{\mathbf{i}}^{(\bm{n})}:=F_{\mathbf{i},n_{1}}\otimes\cdots\otimes F_{\mathbf{i},n_{d}}\in\mathbb{C}_{\mathbf{i}}^{N_{\bm{n}}\times N_{\bm{n}}}$. Given the block sizes $s,t\in\mathbb{N}$ and since, for the time being, we define the DFT in the slice $\mathbb{C}_{\mathbf{i}}$, we use the following compact notation
$U_{L}:=F_{\mathbf{i}}^{(\bm{n})}\otimes I_{s},\qquad U_{R}:=F_{\mathbf{i}}^{(\bm{n})}\otimes I_{t}.$
In this section, most of the time and without loss of generality, we work with right polynomials and right generating functions. Given a right-polynomial, in the following we introduce $d$-level block circulants:
Multilevel block quaternion circulants generated by trigonometric polynomials
We start by recalling the canonical construction of multilevel block circulant matrices in the quaternion setting. First, given $\bm{n}=(n_{1},\ldots,n_{d})\in\mathbb{N}^{d}$ and the multi-index set $\Lambda_{\bm{n}}$, the elementary periodic multilevel shifts are defined by
$P_{\bm{n}}^{(\bm{\rho})}:=P_{n_{1}}^{(\rho_{1})}\otimes\cdots\otimes P_{n_{d}}^{(\rho_{d})},\qquad[P_{n_{k}}^{(\rho_{k})}]_{ij}:=\mathbf{1}_{\{i\equiv j+\rho_{k}\ (\mathrm{mod}\ n_{k})\}}.$
We now define multilevel block circulant matrices generated by matrix-valued right-polynomials. Most of the time, we use the definition based on periodic shifts, due to its algebraic convenience in our derivations, but we also recall the entrywise definition.
Definition 4.17 ($d$-level $s\times t$ quaternion block-circulant generated by a right polynomial).
Given a right-polynomial $p_{\mathrm{R}}$ with block coefficients $\{p_{\mathrm{R}}(\bm{\rho})\}_{\bm{\rho}\in\mathcal{F}}\subset\mathbb{H}^{s\times t}$, we define the associated $d$-level $s\times t$ block-circulant matrix as
$C_{\bm{n}}\bigl(p_{\mathrm{R}}\bigr):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(P_{\bm{n}}^{(\bm{\rho})}\otimes p_{\mathrm{R}}(\bm{\rho})\bigr)\ \in\ \mathbb{H}^{(N_{\bm{n}}s)\times(N_{\bm{n}}t)},\qquad N_{\bm{n}}=\prod_{\ell=1}^{d}n_{\ell}.$
Equivalently, we can describe the entries explicitly
$\bigl[C_{\bm{n}}\bigl(p_{\mathrm{R}}\bigr)\bigr]_{\alpha,\beta}=\sum_{\bm{\rho}\in\mathcal{F}}p_{\mathrm{R}}(\bm{\rho})\,\mathbf{1}_{\{\alpha\equiv\beta+\bm{\rho}\ (\mathrm{mod}\ \bm{n})\}},$
and let its degree $\deg p_{\mathrm{R}}:=\max_{\bm{\rho}\in\mathcal{F}}\|\bm{\rho}\|_{\infty}$.
To find a suitable block-diagonalization of multilevel block-circulant matrices, we follow the Pan-Ng splitting approach verbatim, in particular the results found in [41]. The approach in the multilevel and block setting is completely analogous. Further, this approach is purely quaternion algebraic and is suitable for our a.c.s  approximation formalism. We note that a full-block diagonalization of unilevel-block circulants and a Fourier-based interpretation of diagonal blocks is also found by Zheng and Ni in [65] by moving the problem from quaternion to non-associative octonions.
First, write each polynomial coefficient in Cartesian form $p_{\mathrm{R}}(\bm{\rho})=p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}+p_{\mathrm{R}}(\bm{\rho})^{(2)}\mathbf{j}+p_{\mathrm{R}}(\bm{\rho})^{(3)}(\mathbf{i}\mathbf{j})$ with $p_{\mathrm{R}}(\bm{\rho})^{(\ell)}\in\mathbb{R}^{s\times t}$. By expanding the shift representation, we obtain
$C_{\bm{n}}\bigl(p_{\mathrm{R}}\bigr)$ $=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(P_{\bm{n}}^{(\bm{\rho})}\otimes p_{\mathrm{R}}(\bm{\rho})^{(0)}\bigr)\;+\;\Bigl(\sum_{\bm{\rho}\in\mathcal{F}}P_{\bm{n}}^{(\bm{\rho})}\otimes p_{\mathrm{R}}(\bm{\rho})^{(1)}\Bigr)\,\mathbf{i}$
$\quad+\;\Bigl(\sum_{\bm{\rho}\in\mathcal{F}}P_{\bm{n}}^{(\bm{\rho})}\otimes p_{\mathrm{R}}(\bm{\rho})^{(2)}\Bigr)\,\mathbf{j}\;+\;\Bigl(\sum_{\bm{\rho}\in\mathcal{F}}P_{\bm{n}}^{(\bm{\rho})}\otimes p_{\mathrm{R}}(\bm{\rho})^{(3)}\Bigr)\,(\mathbf{i}\mathbf{j}).$
In particular, we observe that $C_{\bm{n}}\bigl(p_{\mathrm{R}}\bigr)$ is a right-linear combination of four real multilevel block-circulants. The idea behind our target diagonalization is the observation that applying linearly the QDFT to the circulant components in the complex slice $\mathbb{C}_{\mathbf{i}}$ and to the orthogonal circulant components produces different sparsity patterns, and when we sum all of these the sparsity results in a multilevel version of the “X shape” decomposition found in [41].
Next, we need the following elementary algebraic identities of QDFTs.
Lemma 4.18 (QDFT identities, 1D and multilevel).
Fix $\mathbf{i}\in\operatorname{Im}\mathbb{H}$ and identify $\mathbb{C}_{\mathbf{i}}\cong\mathbb{C}$. For $n\in\mathbb{N}$ let $F_{\mathbf{i},n}\in\mathbb{C}_{\mathbf{i}}^{n\times n}$ be $F_{\mathbf{i},n}]_{uv}=n^{-1/2}\exp(-2\pi\mathbf{i}\,uv/n)$, $u,v=0,\dots,n-1$. Then
$F_{\mathbf{i},n}^{*}F_{\mathbf{i},n}=I_{n},\qquad F_{\mathbf{i},n}^{T}=F_{\mathbf{i},n},\qquad(F_{\mathbf{i},n})^{2}=A_{n},$
where $A_{n}$ is the reversal permutation (first/last, second/second-last, …). For $\bm{n}=(n_{1},\dots,n_{d})$ and $F_{\mathbf{i}}^{(\bm{n})}:=\bigotimes_{\ell=1}^{d}F_{\mathbf{i},n_{\ell}}$,
$\bigl(F_{\mathbf{i}}^{(\bm{n})}\bigr)^{*}F_{\mathbf{i}}^{(\bm{n})}=I_{N},\quad\bigl(F_{\mathbf{i}}^{(\bm{n})}\bigr)^{T}=F_{\mathbf{i}}^{(\bm{n})},\quad\bigl(F_{\mathbf{i}}^{(\bm{n})}\bigr)^{2}=A_{\bm{n}}:=\bigotimes_{\ell=1}^{d}A_{n_{\ell}},$
with $N=\prod_{\ell=1}^{d}n_{\ell}$.
As a consequence, we also obtain QDFT flips for multilevel block transforms.
Lemma 4.19.
For any $X\in\mathbb{C}_{\mathbf{i}}^{m\times n}$ one has
$F_{\mathbf{i}}^{(\bm{n})}\,\mathbf{j}\,\bigl(F_{\mathbf{i}}^{(\bm{n})}\bigr)^{*}=A_{\bm{n}}\,\mathbf{j},\qquad F_{\mathbf{i}}^{(\bm{n})}\,(\mathbf{i}\mathbf{j})\,\bigl(F_{\mathbf{i}}^{(\bm{n})}\bigr)^{*}=A_{\bm{n}}\,(\mathbf{i}\mathbf{j}).$ (4.22)
The DFT-canonical form of multilevel block-circulant matrices is folklore in numerical linear algebra.
Lemma 4.20 (Real multilevel block-circulants canonical form).
Let $H:\Lambda_{\bm{n}}\to\mathbb{R}^{s\times t}$ and $S=\sum_{\bm{\rho}\in\Lambda_{\bm{n}}}P_{\bm{n}}^{(\bm{\rho})}\otimes H(\bm{\rho})$ be a real $d$-level $s\times t$ block-circulant. Then
$U_{L}\,S\,U_{R}^{*}=\operatorname{Bdiag}\bigl(\widehat{H}(\bm{k})\bigr)_{\bm{k}\in\Lambda_{\bm{n}}},\quad\widehat{H}(\bm{k})=\sum_{\bm{\rho}\in\Lambda_{\bm{n}}}H(\bm{\rho})\,e^{-2\pi\mathbf{i}\sum_{\ell=1}^{d}k_{\ell}\rho_{\ell}/n_{\ell}}.$
To obtain a block-diagonal sparsity pattern, we need to use permutations. The following ordering description is the multilevel generalization of the one used by [41]:
Definition 4.21.
Let $d\in\mathbb{N}$, $\bm{n}=(n_{1},\ldots,n_{d})\in\mathbb{N}^{d}$ and $\Lambda_{\bm{n}}:=\prod_{\ell=1}^{d}\{0,1,\ldots,n_{\ell}-1\}$ the standard Cartesian index set. If $\bm{k}=(k_{1},\ldots,k_{d})\in\Lambda_{\bm{n}}$, then we write $\bm{k}\equiv((-k_{1})\bmod n_{1},\ldots,(-k_{d})\bmod n_{d})$. Define the fixed-point set
$\mathrm{Fix}:=\{\bm{k}\in\Lambda_{\bm{n}}:\ \bm{k}\equiv-\bm{k}\ (\mathrm{mod}\ \bm{n})\}=\bigl\{\bm{k}\in\Lambda_{\bm{n}}:\ 2k_{\ell}\equiv 0\ (\mathrm{mod}\ n_{\ell})\ \forall\,\ell\bigr\}.$
Definition 4.22.
We denote the standard lexicographic order on $\Lambda_{\bm{n}}$ by $\prec$, and set
$\mathcal{K}:=\{\bm{k}\in\Lambda_{\bm{n}}\setminus\mathrm{Fix}:\ \bm{k}\prec-\bm{k}\}.$
Then we define the ordered index set
$\mathcal{L}=\Bigl[\ \underbrace{\bm{k}^{(1)},\ldots,\bm{k}^{(|\mathrm{Fix}|)}}_{\bm{k}^{(j)}\in\mathrm{Fix}}\ \ \big|\ \ \underbrace{\bm{k}_{1},-\bm{k}_{1},\ \bm{k}_{2},-\bm{k}_{2},\ \ldots}_{\bm{k}_{j}\in\mathcal{K}\\ }\Bigr],$
where $\bm{k}^{(i)}$ and $\bm{k}_{j}$ are both ordered in lexicographic order.
Let $P$ be the permutation matrix that maps the canonical basis according to the order law given by $\mathcal{L}$. For the multilevel $s\times t$ block case, we define $\Pi_{L}:=P\otimes I_{s}$ and $\Pi_{R}:=P\otimes I_{t}$.
By construction, we have the following identities:
Lemma 4.23.
Let $A_{\bm{n}}=\bigotimes_{\ell=1}^{d}A_{n_{\ell}}$ be the multilevel exchange permutation as in Lemma 4.18. With $P$ from Definition 4.22 and $\mathcal{K}$ as above,
$P\,A_{\bm{n}}\,P^{*}=I_{|\mathrm{Fix}|}\ \oplus\ \bigoplus_{\bm{k}\in\mathcal{K}}\begin{bmatrix}0&1\\[2.0pt] 1&0\end{bmatrix}$.
Consequently,
$\Pi_{R}\,(A_{\bm{n}}\otimes I_{t})\,\Pi_{R}^{*}=I_{|\mathrm{Fix}|}\otimes I_{t}\ \oplus\ \bigoplus_{\bm{k}\in\mathcal{K}}\begin{bmatrix}0&I_{t}\\[2.0pt] I_{t}&0\end{bmatrix}$.
Given these premises, we can derive the canonical QDFT decomposition for multilevel block-circulant matrices:
Proposition 4.24 (Canonical QDFT, multilevel rectangular block case).
Let $U_{L}:=F_{\mathbf{i}}^{(\bm{n})}\otimes I_{s}$ and $U_{R}:=F_{\mathbf{i}}^{(\bm{n})}\otimes I_{t}$. Define, for $\ell=0,1,2,3$,
$\Lambda_{\ell}:=\operatorname{Bdiag}\bigl(\widehat{S}_{\ell}(\bm{k})\bigr)_{\bm{k}\in\Lambda_{\bm{n}}},$
where
$\widehat{S}_{\ell}(\bm{k})=\sum_{\bm{\rho}\in\mathcal{F}}p_{\mathrm{R}}(\bm{\rho})^{(\ell)}\,e^{-2\pi\mathbf{i}\sum_{\ell^{\prime}=1}^{d}k_{\ell^{\prime}}\rho_{\ell^{\prime}}/n_{\ell^{\prime}}}.$
Then
$U_{L}\,C_{\bm{n}}\bigl(p_{\mathrm{R}}\bigr)\,U_{R}^{*}=\Lambda_{0}+\Lambda_{1}\,\mathbf{i}+\Lambda_{2}\,(A_{\bm{n}}\otimes I_{t})\,\mathbf{j}+\Lambda_{3}\,(A_{\bm{n}}\otimes I_{t})\,(\mathbf{i}\mathbf{j}).$ (4.23)
Let $P$ be the paired permutation of Definition 4.22, and set $\Pi_{L}:=P\otimes I_{s}$, $\Pi_{R}:=P\otimes I_{t}$. After conjugation by $\Pi_{L},\Pi_{R}$ one has the block direct sum
$\Pi_{L}U_{L}\,C_{\bm{n}}\bigl(p_{\mathrm{R}}\bigr)\,U_{R}^{*}\Pi_{R}^{*}=\bigoplus_{\bm{k}\in\mathrm{Fix}}\bigl(D_{1}(\bm{k}){+D_{2}(\bm{k})}\bigr)\ \ \oplus\ \bigoplus_{\{\bm{k},-\bm{k}\}}\begin{bmatrix}D_{1}(\bm{k})&D_{2}(\bm{k})\\[2.0pt] D_{2}(-\bm{k})&D_{1}(-\bm{k})\end{bmatrix}$,
where
$D_{1}(\bm{k}):=\widehat{S}_{0}(\bm{k})+\widehat{S}_{1}(\bm{k})\,\mathbf{i},\qquad D_{2}(\bm{k}):=\bigl(\widehat{S}_{2}(\bm{k})+\widehat{S}_{3}(\bm{k})\,\mathbf{i}\bigr)\,\mathbf{j}.$
Proof.
Write each coefficient of $p_{\mathrm{R}}$ in Cartesian form
$p_{\mathrm{R}}(\bm{\rho})=p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}+p_{\mathrm{R}}(\bm{\rho})^{(2)}\mathbf{j}+p_{\mathrm{R}}(\bm{\rho})^{(3)}(\mathbf{i}\mathbf{j}),$
with real blocks $p_{\mathrm{R}}(\bm{\rho})^{(\ell)}\in\mathbb{R}^{s\times t}$. Expanding the shift representation,
$C_{\bm{n}}\bigl(p_{\mathrm{R}}\bigr)$ $=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(P_{\bm{n}}^{(\bm{\rho})}\otimes p_{\mathrm{R}}(\bm{\rho})^{(0)}\bigr)+\Bigl(\sum_{\bm{\rho}\in\mathcal{F}}P_{\bm{n}}^{(\bm{\rho})}\otimes p_{\mathrm{R}}(\bm{\rho})^{(1)}\Bigr)\,\mathbf{i}$
$\qquad+\Bigl(\sum_{\bm{\rho}\in\mathcal{F}}P_{\bm{n}}^{(\bm{\rho})}\otimes p_{\mathrm{R}}(\bm{\rho})^{(2)}\Bigr)\,\mathbf{j}+\Bigl(\sum_{\bm{\rho}\in\mathcal{F}}P_{\bm{n}}^{(\bm{\rho})}\otimes p_{\mathrm{R}}(\bm{\rho})^{(3)}\Bigr)\,(\mathbf{i}\mathbf{j}).$
For $\ell=0,1$, Lemma 4.20 applied componentwise gives
$U_{L}\Bigl(\sum_{\bm{\rho}}P_{\bm{n}}^{(\bm{\rho})}\otimes p_{\mathrm{R}}(\bm{\rho})^{(\ell)}\Bigr)U_{R}^{*}=\operatorname{Bdiag}\bigl(\widehat{S}_{\ell}(\bm{k})\bigr)_{\bm{k}\in\Lambda_{\bm{n}}}=:\Lambda_{\ell}.$
For the $\mathbf{j}$-component we use that $\mathbf{j}\alpha=\overline{\alpha}\,\mathbf{j}$ for $\alpha\in\mathbb{C}_{\mathbf{i}}$ and the QDFT flip $F_{\mathbf{i}}^{(\bm{n})}\mathbf{j}(F_{\mathbf{i}}^{(\bm{n})})^{*}=A_{\bm{n}}\mathbf{j}$ (Lemma 4.19). As a consequence
$\ U_{L}\Bigl(\sum_{\bm{\rho}}P_{\bm{n}}^{(\bm{\rho})}\otimes p_{\mathrm{R}}(\bm{\rho})^{(2)}\Bigr)\mathbf{j}\,U_{R}^{*}$
$=\bigl[U_{L}\Bigl(\sum_{\bm{\rho}}P_{\bm{n}}^{(\bm{\rho})}\otimes p_{\mathrm{R}}(\bm{\rho})^{(2)}\Bigr)U_{R}^{*}\bigr]\cdot\bigl(U_{R}\mathbf{j}\,U_{R}^{*}\bigr)$
$=\Lambda_{2}\cdot\bigl((F_{\mathbf{i}}^{(\bm{n})}\otimes I_{t})\,\mathbf{j}\,(F_{\mathbf{i}}^{(\bm{n})}\otimes I_{t})^{*}\bigr)$
$=\Lambda_{2}\cdot\bigl((F_{\mathbf{i}}^{(\bm{n})}\mathbf{j}(F_{\mathbf{i}}^{(\bm{n})})^{*})\otimes I_{t}\bigr)=\Lambda_{2}\,(A_{\bm{n}}\otimes I_{t})\,\mathbf{j}.$
The case $\ell=3$ is identical, and this gives us (4.23). Consider now the permutations $P$ of Definition 4.22 and set $\Pi_{L}:=P\otimes I_{s}$ and $\Pi_{R}:=P\otimes I_{t}$. Define
$\widetilde{\Lambda}_{\ell}:=\Pi_{L}\Lambda_{\ell}\Pi_{R}^{*}\qquad(\ell=0,1,2,3).$
if we conjugate (4.23) by $\Pi_{L},\Pi_{R}$, we obtain
$\Pi_{L}U_{L}\,C_{\bm{n}}(p_{\mathrm{R}})\,U_{R}^{*}\Pi_{R}^{*}=\widetilde{\Lambda}_{0}+\widetilde{\Lambda}_{1}\,\mathbf{i}+\Pi_{L}\Lambda_{2}(A_{\bm{n}}\otimes I_{t})\Pi_{R}^{*}\,\mathbf{j}+\Pi_{L}\Lambda_{3}(A_{\bm{n}}\otimes I_{t})\Pi_{R}^{*}\,(\mathbf{i}\mathbf{j}).$
Now, insert the identity $\Pi_{R}^{*}\Pi_{R}=I$ between $\Lambda_{2}$ and $(A_{\bm{n}}\otimes I_{t})$ so that
$\Pi_{L}\Lambda_{2}(A_{\bm{n}}\otimes I_{t})\Pi_{R}^{*}=(\Pi_{L}\Lambda_{2}\Pi_{R}^{*})\cdot\bigl(\Pi_{R}(A_{\bm{n}}\otimes I_{t})\Pi_{R}^{*}\bigr),$
and in a similar way, we do the same for $\Lambda_{3}$. By using the identity in block form of Lemma 4.23, we have
$Z:=\Pi_{R}(A_{\bm{n}}\otimes I_{t})\Pi_{R}^{*}=(P\otimes I_{t})(A_{\bm{n}}\otimes I_{t})(P\otimes I_{t})^{*}=I_{|\mathrm{Fix}|}\otimes I_{t}\ \oplus\ \bigoplus_{\bm{k}\in\mathcal{K}}\begin{bmatrix}0&I_{t}\\[2.0pt] I_{t}&0\end{bmatrix}$.
Therefore, we infer
$\Pi_{L}U_{L}\,C_{\bm{n}}(p_{\mathrm{R}})\,U_{R}^{*}\Pi_{R}^{*}=\widetilde{\Lambda}_{0}+\widetilde{\Lambda}_{1}\,\mathbf{i}+\bigl(\widetilde{\Lambda}_{2}Z\bigr)\,\mathbf{j}+\bigl(\widetilde{\Lambda}_{3}Z\bigr)\,(\mathbf{i}\mathbf{j}).$ (4.24)
The relation obtained in (4.24) reads as follows.
  • If $\bm{k}\in\mathrm{Fix}$, we have $Z=I_{t}$, and the block diagonals we obtain are equal to
    $D_{1}(\bm{k}){+D_{2}(\bm{k})}:=\widehat{S}_{0}(\bm{k})+\widehat{S}_{1}(\bm{k})\,\mathbf{i}+{\bigl(\widehat{S}_{2}(\bm{k})+\widehat{S}_{3}(\bm{k})\,\mathbf{i}\bigr)\,\mathbf{j}}.$
  • Instead, for each unordered pair $\{\bm{k},-\bm{k}\}$, the factor $Z$ is the $t$-block $2\times 2$ exchange matrix, and the corresponding blocks we obtain are
    $\begin{bmatrix}D_{1}(\bm{k})&D_{2}(\bm{k})\\[2.0pt] D_{2}(-\bm{k})&D_{1}(-\bm{k})\end{bmatrix},\qquad D_{2}(\bm{k}):=\bigl(\widehat{S}_{2}(\bm{k})+\widehat{S}_{3}(\bm{k})\,\mathbf{i}\bigr)\,\mathbf{j}.$
Collecting all the sums in the block-diagonal order given by $P$ gives exactly our claim. ∎
4.6.1 The spectrum of multilevel block-circulant matrices and the a.c.s. circulant approximation of Toeplitz sequences
In the current section, we discuss the details of the canonical spectrum and singular values of multilevel block circulant matrices. Then we use these details to find the asymptotic spectral/singular value distribution of the corresponding matrix sequence and prove that every quaternion Toeplitz sequence has a multilevel right block-circulant a.c.s.
Let $p_{\mathrm{R}}(\bm{\theta})=\sum_{\bm{\rho}\in\mathcal{F}}p_{\mathrm{R}}(\bm{\rho})\,e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle}$ be a $s\times t$ trigonometric polynomial with finite $\mathcal{F}\subset\mathbb{Z}^{d}$, Construct the multilevel block-circulant $C_{\bm{n}}\bigl(p_{\mathrm{R}}\bigr)$ generated by $\{p_{\mathrm{R}}(\bm{\rho})\}_{\bm{\rho}\in\mathcal{F}}$ as in Definition 4.17 and consider $U_{R}$, $U_{L}$, and $A_{\bm{n}}$ used in the preceding proposition (Proposition 4.24). We need to highlight the fiber property of the block diagonal we obtain with the factorization in Proposition 4.24, meaning that their symplectic embeddings are uniform samples of the spectral and singular value symbol of the corresponding Circulant sequence $\{C_{\bm{n}}\bigl(p_{\mathrm{R}}\bigr)\}_{\bm{n}}$.
Consider the uniform sample set $\Theta$ in $\mathbb{T}^{p}$ with elements $\bm{\theta}_{\bm{k}}=2\pi\bigl(\bm{k}/\bm{n})$ for $\bm{k}\in\Lambda_{\bm{n}}$ and note that, by periodicity $\bm{\theta}_{\bm{k}}=\bm{\theta}_{-\bm{k}}\ (\mathrm{mod}\ 2\pi)$.
Now, write the Cartesian split $p_{\mathrm{R}}(\bm{\rho})=p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}+p_{\mathrm{R}}(\bm{\rho})^{(2)}\mathbf{j}+p_{\mathrm{R}}(\bm{\rho})^{(3)}(\mathbf{i}\mathbf{j})$ with $p_{\mathrm{R}}(\bm{\rho})^{(\ell)}\in\mathbb{R}^{s\times t}$, and define
$Z(\bm{\theta}):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},\qquad W(\bm{\theta}):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(2)}+p_{\mathrm{R}}(\bm{\rho})^{(3)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle}.$ (4.25)
For our aims, we need the symplectic embedding in block form $\Phi_{1}:\mathbb{H}^{s\times t}\to\mathbb{C}^{(2s)\times(2t)}$.
To obtain the samplings of canonical eigenvalues and singular values of circulant matrices, we make use of the following claim, which is a consequence of Proposition 4.24.
Proposition 4.25.
Let $p_{\mathrm{R}}(\bm{\theta})=\sum_{\bm{\rho}\in\mathcal{F}}p_{\mathrm{R}}(\bm{\rho})\,e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle}$ be an $s\times t$ trigonometric polynomial with finite $\mathcal{F}\subset\mathbb{Z}^{d}$ and coefficients $p_{\mathrm{R}}(\bm{\rho})\in\mathbb{H}^{s\times t}$ and consider
$Z(\bm{\theta}):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},\qquad W(\bm{\theta}):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(2)}+p_{\mathrm{R}}(\bm{\rho})^{(3)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},$
defined in (4.25). Let $C_{\bm{n}}(p_{\mathrm{R}})$ be the associated $d$-level $s\times t$ block-circulant.
Then, from the QDFT canonical form of $C_{\bm{n}}(p_{\mathrm{R}})$ written in the notation
$\Pi_{L}\,U_{L}\,C_{\bm{n}}(p_{\mathrm{R}})\,U_{R}^{*}\,\Pi_{R}^{*}=\bigoplus_{\bm{k}\in\mathrm{Fix}}\bigl(Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)\ \ \oplus\ \bigoplus_{\{\bm{k},-\bm{k}\}\subset\Lambda_{\bm{n}}\setminus\mathrm{Fix}}\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}$, (4.26)
we obtain the embedding of the block diagonals:
$\Phi\!\left(\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}\right)=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&0&0&W(\bm{\theta}_{\bm{k}})\\[2.0pt] 0&\overline{Z(\bm{\theta}_{\bm{k}})}&-\overline{W(\bm{\theta}_{\bm{k}})}&0\\[2.0pt] 0&W(\bm{\theta}_{-\bm{k}})&Z(\bm{\theta}_{-\bm{k}})&0\\[2.0pt] -\overline{W(\bm{\theta}_{-\bm{k}})}&0&0&\overline{Z(\bm{\theta}_{-\bm{k}})}\end{bmatrix}$, (4.27)
when $\bm{k}\in\Lambda_{\bm{n}}$, while for $\bm{k}\in\mathrm{Fix}$ we have
$\Phi\bigl(Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\\[2.0pt] -\overline{W(\bm{\theta}_{\bm{k}})}&\overline{Z(\bm{\theta}_{\bm{k}})}\end{bmatrix}$. (4.28)
Proof.
First, we recall the notation of the canonical form in (4.26). Indeed, by Proposition 4.24, we deduce the canonical QDFT form
$\Pi_{L}\,U_{L}\,C_{\bm{n}}\bigl(p_{\mathrm{R}}\bigr)\,U_{R}^{*}\,\Pi_{R}^{*}=\bigoplus_{\bm{k}\in\mathrm{Fix}}\!\Bigl(\widehat{S}_{0}(\bm{k})+\widehat{S}_{1}(\bm{k})\mathbf{i}+\bigl(\widehat{S}_{2}(\bm{k})+\widehat{S}_{3}(\bm{k})\mathbf{i}\bigr)\mathbf{j}\Bigr)$
$\oplus\ \bigoplus_{\{\bm{k},-\bm{k}\}}\!\begin{bmatrix}\widehat{S}_{0}(\bm{k})+\widehat{S}_{1}(\bm{k})\mathbf{i}&\bigl(\widehat{S}_{2}(\bm{k})+\widehat{S}_{3}(\bm{k})\mathbf{i}\bigr)\mathbf{j}\\[2.0pt] \bigl(\widehat{S}_{2}(-\bm{k})+\widehat{S}_{3}(-\bm{k})\mathbf{i}\bigr)\mathbf{j}&\widehat{S}_{0}(-\bm{k})+\widehat{S}_{1}(-\bm{k})\mathbf{i}\end{bmatrix}$.
Now, by the definitions of $Z$ and $W$ we have
$Z(\bm{\theta}_{\bm{k}})=\widehat{S}_{0}(\bm{k})+\widehat{S}_{1}(\bm{k})\,\mathbf{i},\qquad W(\bm{\theta}_{\bm{k}})=\widehat{S}_{2}(\bm{k})+\widehat{S}_{3}(\bm{k})\,\mathbf{i}.$
Thus the fixed blocks are indeed $Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}$ and the paired blocks are
$\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}$,
which confirms (4.26).
For the complex embedding claim, we have to recall the basic identities for every $X\in\mathbb{C}_{\mathbf{i}}^{s\times t}$
$\Phi(X)=\begin{bmatrix}X&0\\[2.0pt] 0&\overline{X}\end{bmatrix},\qquad\Phi(X\mathbf{j})=\begin{bmatrix}0&X\\[2.0pt] -\overline{X}&0\end{bmatrix}$,
together with the $\mathbb{R}$-linearity of $\Phi$ and its blockwise action on real $2\times 2$ partitions. For a fixed index $\bm{k}\in\mathrm{Fix}$ we obtain
$\Phi\bigl(Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)=\Phi\bigl(Z(\bm{\theta}_{\bm{k}})\bigr)+\Phi\bigl(W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\\[2.0pt] -\overline{W(\bm{\theta}_{\bm{k}})}&\overline{Z(\bm{\theta}_{\bm{k}})}\end{bmatrix}$,
which is (4.28). For the paired blocks, split them as sums of their diagonal and off-diagonal parts,
$\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&0\\[2.0pt] 0&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}\;+\;\begin{bmatrix}0&W(\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(\bm{\theta}_{-\bm{k}})\mathbf{j}&0\end{bmatrix}$.
Applying $\Phi$ to each part and summing, using the two identities above, yields exactly the $4\times 4$ complex block in (4.27). This completes the proof. ∎
As a corollary, we deduce the asymptotic singular value and canonical spectral distribution of multilevel $s\times t$ circulant matrix sequences as reported in the subsequent result.
Corollary 4.26 (Asymptotic spectral/singular value distributions for right-polynomial quaternion circulants).
Fix $\mathbf{i}\in\operatorname{Im}\mathbb{H}$ and let
$p_{\mathrm{R}}(\bm{\theta})=\sum_{\bm{\rho}\in\mathcal{F}}p_{\mathrm{R}}(\bm{\rho})\,e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},\qquad\mathcal{F}\subset\mathbb{Z}^{d}\ \text{finite},\ \ p_{\mathrm{R}}(\bm{\rho})\in\mathbb{H}^{s\times t}.$
Write $p_{\mathrm{R}}(\bm{\rho})=p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}+p_{\mathrm{R}}(\bm{\rho})^{(2)}\mathbf{j}+p_{\mathrm{R}}(\bm{\rho})^{(3)}(\mathbf{i}\mathbf{j})$ with $p_{\mathrm{R}}(\bm{\rho})^{(\ell)}\in\mathbb{R}^{s\times t}$ and define
$Z(\bm{\theta})=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},\qquad W(\bm{\theta})=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(2)}+p_{\mathrm{R}}(\bm{\rho})^{(3)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},$
$G_{\mathrm{R}}\bigl(p_{\mathrm{R}}(\bm{\theta})\bigr):=\begin{bmatrix}Z(\bm{\theta})&W(\bm{\theta})\\[2.0pt] -\overline{W(-\bm{\theta})}&\overline{Z(-\bm{\theta})}\end{bmatrix}$.
Let $C_{\bm{n}}(p_{\mathrm{R}})$ be the corresponding $d$-level $s\times t$ block-circulant. Then
$\{C_{\bm{n}}(p_{\mathrm{R}})\}_{\bm{n}}\ \sim_{\sigma}\ G_{\mathrm{R}}.$
If $s=t$, then also
$\{C_{\bm{n}}(p_{\mathrm{R}})\}_{\bm{n}}\ \sim_{\lambda}\ G_{\mathrm{R}}.$
Proof.
Let $\bm{\theta}_{\bm{k}}=2\pi(\bm{k}/\bm{n})$ for $\bm{k}\in\Lambda_{\bm{n}}$ and $\mathrm{Fix}=\{\bm{k}:2\bm{k}\equiv 0\pmod{\bm{n}}\}$ the usual fixed point set. By Proposition 4.25 (see (4.26)), after the application of the QDFT and the permutations, we block-diagonalize $C_{\bm{n}}(p_{\mathrm{R}})$ as
$\bigl(Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)\quad(\bm{k}\in\mathrm{Fix}),\qquad\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}\quad(\{\bm{k},-\bm{k}\}),$
and after the symplectic embedding we have the $4\times 4$ complex block in (4.27) for pairs and the $2\times 2$ block in (4.28) for fixed indices. Now, if we apply for both rows and columns a $s\times t$ block permutation with the law $(1,4,3,2)$ to the $4\times 4$ blocks (4.27), we transform them into
$\operatorname{diag}\!\Bigl(G_{\mathrm{R}}\bigl(p_{\mathrm{R}}(\bm{\theta}_{\bm{k}})\bigr),\ G_{\mathrm{R}}\bigl(p_{\mathrm{R}}(\bm{\theta}_{-\bm{k}})\bigr)\Bigr),$
while for $\bm{k}\in\mathrm{Fix}$ we have the periodic congruence $\bm{\theta}_{\bm{k}}\equiv\bm{\theta}_{\bm{k}}\ (\mathrm{mod}\ 2\pi)$, so
$\Phi\bigl(Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\\[2.0pt] -\overline{W(\bm{\theta}_{\bm{k}})}&\overline{Z(\bm{\theta}_{\bm{k}})}\end{bmatrix}=G_{\mathrm{R}}\bigl(p_{\mathrm{R}}(\bm{\theta}_{\bm{k}})\bigr).$
Hence, for every $\bm{n}$, we deduce
$\sigma\!\bigl(\Phi(C_{\bm{n}}(p_{\mathrm{R}}))\bigr)=\bigsqcup_{\bm{k}\in\Lambda_{\bm{n}}}\sigma\!\bigl(G_{\mathrm{R}}(p_{\mathrm{R}}(\bm{\theta}_{\bm{k}}))\bigr).$
If $s=t$, the left/right QDFT and the permutation used above coincide so the singular value result is obtained by unitary similarity, and thus we can give the same conclusion for eigenvalues:
$\lambda\!\bigl(\Phi(C_{\bm{n}}(p_{\mathrm{R}}))\bigr)=\bigsqcup_{\bm{k}\in\Lambda_{\bm{n}}}\lambda\!\bigl(G_{\mathrm{R}}(p_{\mathrm{R}}(\bm{\theta}_{\bm{k}}))\bigr).$
As a consequence, note that for any continuous compactly supported test function $\psi$, the Riemann sums on $\{\bm{\theta}_{\bm{k}}\}_{\bm{k}\in\Lambda_{\bm{n}}}$ of the left side of (3.1) and (3.3) applied to $\{C_{\bm{n}}(p_{\mathrm{R}})\}_{\bm{n}}$ are a converging trapezoidal rule that yields $\{C_{\bm{n}}(p_{\mathrm{R}})\}_{\bm{n}}\sim_{\sigma}G_{\mathrm{R}}$ and, when $s=t$, also $\{C_{\bm{n}}(p_{\mathrm{R}})\}_{\bm{n}}\sim_{\lambda}G_{\mathrm{R}}$. ∎
Theorem 4.27 (Right Toeplitz admits an a.c.s. of circulants).
Let $d\in\mathbb{N}$, $s,t\in\mathbb{N}$, and $F\in L^{1}(\mathbb{T}^{d};\mathbb{H}^{s\times t})$ and consider the $d$-level Toeplitz sequence $\{\,T_{\bm{n}}^{(\tau)}(F)\,\}_{\bm{n}}$} as in (4.2). Then there exists a double sequence of right circulants
${B_{\bm{n},m}\ :=\ \sum_{\bm{m}\in\mathcal{S}_{m}}\bigl(P_{\bm{n}}^{(\bm{m})}\otimes\widehat{F}^{(\tau)}({\bm{m}})\bigr)},\qquad\mathcal{S}_{m}\subset\mathbb{Z}^{d}\ \text{finite},\ \ \mathcal{S}_{m}\uparrow\mathbb{Z}^{d},$
such that $\{\{\,B_{\bm{n},m}\,\}_{\bm{n}}\,\}_{\bm{m}}$ is an a.c.s. for $\{T_{\bm{n}}^{(\tau)}(F)\}_{\bm{n}}$ in the sense of Definition 3.8.
Proof.
First, by using Proposition 4.1 and by observing that the tranformed symbol (4.3) is still integrable, without loss of generality we can assume that the Toeplitz sequence to approximate is in the right-kernel form. Now, we can reduce the problem to The proof is completely analogous to the complex multilevel case (see [21, Thm. 3.5]). Indeed, we show the claim when the generating function of $\{T_{\bm{n}}^{(R)}(F)\}_{\bm{n}}$ is a trigonometric polynomials and we extend the result to integrable functions using basic trigonometric approximation arguments. In fact, since trigonometric polynomials are dense in $L^{1}(\mathbb{T}^{d})$ by Proposition 2.16 and Fourier Series converge in $L^{1}(\mathbb{T}^{p},\mathbb{H}^{s\times t})$ we can find finite sets of indices $\mathcal{S}_{m}\subset\mathbb{Z}^{d}$ and define
$F_{m}(\bm{\theta})\ =\ \sum_{\bm{m}\in\mathcal{S}_{m}}\widehat{F}^{\mathrm{R}}\!\left(\bm{m}\right)\,e^{-\mathbf{i}\langle\bm{m},\bm{\theta}\rangle}\quad\text{such that}\quad\|F-F_{m}\|_{L^{1}(\mathbb{T}^{d})}\xrightarrow[m\to\infty]{}0.$
Fix $m$ and write $\mathcal{S}_{m}\subset\prod_{\ell=1}^{d}\{-r_{\ell}(m),\dots,r_{\ell}(m)\}$. For all $\bm{n}$ with $n_{\ell}\geq r_{\ell}(m)+1$,
$T_{\bm{n}}^{(R)}(F_{m})-\sum_{\bm{m}\in\mathcal{S}_{m}}\bigl(P_{\bm{n}}^{(\bm{m})}\otimes\widehat{F}^{\mathrm{R}}\!\left(\bm{m}\right)\bigr)=\sum_{\bm{m}\in\mathcal{S}_{m}}\bigl(J_{\bm{n}}^{(\bm{m})}-P_{\bm{n}}^{(\bm{m})}\bigr)\otimes\widehat{F}^{\mathrm{R}}\!\left(\bm{m}\right),$
where $J_{\bm{n}}^{(\bm{m})}$ denotes the (real) Toeplitz shift. For each $\bm{m}=(m_{1},\dots,m_{d})$, it is known that
$\operatorname{rank}\!\bigl(J_{\bm{n}}^{(\bm{m})}-P_{\bm{n}}^{(\bm{m})}\bigr)\ \leq\ \sum_{\ell=1}^{d}2|m_{\ell}|\,\frac{N_{\bm{n}}}{n_{\ell}}.$
Using $\operatorname{rank}(X\otimes Y)\leq\operatorname{rank}(X)\operatorname{rank}(Y)$ and $\operatorname{rank}\widehat{F}^{\mathrm{R}}\!\left(\bm{m}\right)\leq\min\{s,t\}$, with $r_{\bm{n}}:=N_{\bm{n}}\min\{s,t\}$, we obtain
$\frac{\operatorname{rank}_{\mathbb{H}}\!\Bigl(T_{\bm{n}}^{(R)}(F_{m})-B_{\bm{n},m}\Bigr)}{r_{\bm{n}}}\ \leq\ 2\sum_{\ell=1}^{d}\frac{1}{n_{\ell}}\sum_{\bm{m}\in\mathcal{S}_{m}}|m_{\ell}|\ \xrightarrow[\bm{n}\to\infty]{}\ 0\qquad(\text{for each fixed }m).$
By Theorem 4.13 with $p=1$,
$\|T_{\bm{n}}^{(R)}(F)-T_{\bm{n}}^{(R)}(F_{m})\|_{1}\ \leq\ {4}\,\frac{N_{\bm{n}}}{(2\pi)^{d}}\ \|F-F_{m}\|_{L^{1}(\mathbb{T}^{d})}.$
Dividing by $r_{\bm{n}}=N_{\bm{n}}\min\{s,t\}$ gives
$\frac{\|T_{\bm{n}}^{(R)}(F)-T_{\bm{n}}^{(R)}(F_{m})\|_{1}}{r_{\bm{n}}}\ \leq\ \frac{2}{(2\pi)^{d}\min\{s,t\}}\ \|F-F_{m}\|_{L^{1}(\mathbb{T}^{d})}\ :=:\ \omega_{1}(m)\xrightarrow[m\to\infty]{}0.$
Hence, by Definition 3.8 and $\text{ACS}_{\mathbb{H}}5$, $\{\{B_{\bm{n},m}\}_{\bm{n}}:\ m\}$ is an a.c.s. for $\{T_{\bm{n}}^{(R)}(F)\}_{\bm{n}}$. ∎
Now, we report briefly how to compute the singular value and spectral distribution of Toeplitz sequences as in Theorem 4.8 using the classical a.c.s approximation approach with the tools furnished in the current subsection.
a.c.s  based proof of Theorem 4.8.
The claim follows directly from [21, Thm. 3.5], using $\text{ACS}_{\mathbb{H}}1$ and $\text{ACS}_{\mathbb{H}}5$ since
  1. 1.
    $\{T_{\bm{n}}^{(\tau)}(F)\}_{\bm{n}}$ admits a quaternion a.c.s $\{\{B_{\bm{n},m}\}_{\bm{n}}:\ m\}$ composed by circulant matrices as in Theorem 4.27;
  2. 2.
    By construction of Theorem 4.27, $\{B_{\bm{n},m}\}_{n}\sim_{\lambda\,,\,\sigma}F_{m}$, for some trigonometric polynomials $F_{m}$
  3. 3.
    $F_{m}\to F$ in $L^{1}$.
Compare the hypotheses given here to the ones in [21, Thm. 3.5], in which the a.c.s based proof of the spectral distribution of Toeplitz sequences is explained in more detail. ∎
5 Numerical Results
In this section, we present numerical tests focused on showing the results of Theorem 4.8 regarding the singular value distribution of the Toeplitz quaternion sequences and the spectral distribution in the Hermitian setting empirically. The experimental setting is standard in asymptotic spectral analysis: indeed we compare the non-decreasing sorted eigenvalues of the Toeplitz matrix against a fine uniform sampling of a monotone non-decreasing rearrangement of the symbol (see [21] for the definition of monotone rearrangement we use). Owing to Theorem 4.8, we expect an evident match between the two plots as the size of the matrix is large. We further remark that, as observed in [7] and [3], this experimental setup coincides with verifying numerically the $L^{1}$ convergence of quantile functions.
More specifically, our experiments focus on showing quantile convergence for twolevel block quaternionic Toeplitz matrices generated by square $2\times 2$ twolevel block generating functions. This choice makes the test general, computational tractable, and spectrally informative, if applied, for example, to dual-channel 2D systems in signal analysis problems. We finally anticipate that the match is quite good even for moderate matrix-sizes, while the statements contained in Theorem 4.8 have asymptotical nature.
Computational Setup
All numerical experiments are performed on a machine that features:
  • CPU: Intel Core i7-1355U
  • RAM: 16 GB
  • OS: Windows 11 (64-bit)
The code was written in Python 3.11 (Anaconda distribution), using the following packages:
  • numpy (linear algebra, FFTs)
  • matplotlib (plots and visualization)
  • quaternion / numpy-quaternion (quaternion arithmetic)
All experiments are executed in a standard CPU environment.
5.1 Continuous symbols
Tests use matrix sizes $(n,n)$ with $n=2,8,32$ for both Hermitian and non-Hermitian cases.
Continuous Hermitian.
We consider the following smooth symmetric function:
$F(\theta_{1},\theta_{2})=\begin{bmatrix}2+\cos\theta_{1}+\cos\theta_{2}&\tfrac{1}{2}\bigl(\sin\theta_{1}\,\mathbf{i}+\sin\theta_{2}\,\mathbf{j}\bigr)+\tfrac{1}{4}\sin(\theta_{1}-\theta_{2})\,\mathbf{k}\\[4.0pt] \tfrac{1}{2}\bigl(\sin\theta_{2}\,\mathbf{j}-\sin\theta_{1}\,\mathbf{i}\bigr)+\tfrac{1}{4}\sin(\theta_{1}-\theta_{2})\,\mathbf{k}&2+\cos\theta_{1}+\cos\theta_{2}\end{bmatrix}$.
Left (L; $S_{L}=\{1,2\},\ S_{R}=\varnothing$)
Refer to caption Refer to caption
Sandwich ($S_{12}$; $S_{L}=\{1\},\ S_{R}=\{2\}$)
Refer to caption Refer to caption
Sandwich ($S_{21}$; $S_{L}=\{2\},\ S_{R}=\{1\}$)
Refer to caption Refer to caption
Right (R; $S_{L}=\varnothing,\ S_{R}=\{1,2\}$)
Refer to caption Refer to caption
[…]
Continuous non-Hermitian.
Consider the following non-symmetric smooth function:
$F(\theta_{1},\theta_{2})=\begin{bmatrix}e^{\mathbf{i}\theta_{1}}+e^{\mathbf{j}\theta_{2}}&(\cos\theta_{2}+\sin\theta_{1})\,\mathbf{k}\\[4.0pt] (\cos\theta_{1}-\sin\theta_{2})\,\mathbf{i}&1+\tfrac{1}{2}\bigl(e^{\mathbf{j}\theta_{1}}-e^{\mathbf{k}\theta_{2}}\bigr)\end{bmatrix}$.
Left (L; $S_{L}=\{1,2\},\ S_{R}=\varnothing$)
Refer to caption Refer to caption
Sandwich ($S_{12}$; $S_{L}=\{1\},\ S_{R}=\{2\}$)
Refer to caption Refer to caption
Sandwich ($S_{21}$; $S_{L}=\{2\},\ S_{R}=\{1\}$)
Refer to caption Refer to caption
Right (R; $S_{L}=\varnothing,\ S_{R}=\{1,2\}$)
Refer to caption Refer to caption
[…]
5.2 $L^{1}$ symbols
Tests use matrix sizes $(n,n)$ with $n=2,8,32$ for both Hermitian and non-Hermitian cases.
In the $L^{1}$ examples below, discontinuities are modeled using the sign function $\operatorname{sgn}(x)$ defined by
$\operatorname{sgn}(x)=\begin{cases}+1,&x>0,\\[2.0pt] 0,&x=0,\\[2.0pt] -1,&x<0.\end{cases}$
$L^{1}$ Hermitian.
A piecewise-constant hermitian symbol:
$F(\theta_{1},\theta_{2})=\begin{bmatrix}2&\operatorname{sgn}(\theta_{1})\,\mathbf{i}+\operatorname{sgn}(\theta_{2})(\mathbf{j}+\mathbf{k})\\[4.0pt] -\operatorname{sgn}(\theta_{1})\,\mathbf{i}+\operatorname{sgn}(\theta_{2})(\mathbf{j}+\mathbf{k})&2\end{bmatrix}$.
Left (L; $S_{L}=\{1,2\},\ S_{R}=\varnothing$)
Refer to caption Refer to caption
Sandwich ($S_{12}$; $S_{L}=\{1\},\ S_{R}=\{2\}$)
Refer to caption Refer to caption
Sandwich ($S_{21}$; $S_{L}=\{2\},\ S_{R}=\{1\}$)
Refer to caption Refer to caption
Right (R; $S_{L}=\varnothing,\ S_{R}=\{1,2\}$)
Refer to caption Refer to caption
[…]
$L^{1}$ non-Hermitian.
As last example, consider the following non-symmetric function:
$F(\theta_{1},\theta_{2})=\begin{bmatrix}e^{\,\mathbf{i}\theta_{1}}+\operatorname{sgn}(\theta_{1})+\mathbf{j}\,\operatorname{sgn}(\theta_{2})&\mathbf{k}\,\operatorname{sgn}\!\bigl(\sin\theta_{1}+\tfrac{1}{4}\cos\theta_{2}\bigr)\\[6.0pt] \mathbf{i}\,\operatorname{sgn}\!\bigl(\cos\theta_{1}-\sin\theta_{2}\bigr)&1+\tfrac{1}{4}\!\left(e^{\,\mathbf{j}\,\operatorname{sgn}(\theta_{1})\pi/2}-e^{\,\mathbf{k}\,\operatorname{sgn}(\theta_{2})\pi/2}\right)\end{bmatrix}$.
Left (L; $S_{L}=\{1,2\},\ S_{R}=\varnothing$)
Refer to caption Refer to caption
Sandwich ($S_{12}$; $S_{L}=\{1\},\ S_{R}=\{2\}$)
Refer to caption Refer to caption
Sandwich ($S_{21}$; $S_{L}=\{2\},\ S_{R}=\{1\}$)
Refer to caption Refer to caption
Right (R; $S_{L}=\varnothing,\ S_{R}=\{1,2\}$)
Refer to caption Refer to caption
[…]
Across all regimes-continuous vs. $L^{1}$, Hermitian vs. non-Hermitian-the empirical quantile plot converges toward the theoretical spectral quantiles of the generating symbols as $\mathbf{n}$ increases, confirming the asymptotic distribution theorems on $\mathbb{T}^{2}$.
As a final remark, we observe that the agreement between the empirical distributions (of eigenvalues and singular values) and the symbols deduced theoretically is remarkably precise even for small sizes of the quaternion matrices. This fact, which is not covered by our theoretical investigations, should be studied carefully, having in mind quantitative estimates of the convergence speed.
6 Conclusions and a plan for future investigations
We have presented an exhaustive treatment of Weyl eigenvalue and singular value distributions for single-axis quaternion block multilevel Toeplitz matrix sequences generated by $s\times t$ quaternion matrix-valued $d$-variate Lebesgue integrable generating functions: this study solves in the positive a large part of the program given in RP 4, items 1, 2, in [52]. Furthermore, we have proved localization results in the quaternion Hermitian setting and we have generalized to the specific quaternion case the notion of approximating class of sequences: in particular, we have shown that block multilevel circulant matrix sequences are good approximation candidates for block multilevel Toeplitz matrix sequences in the a.c.s. sense, to be used for preconditioning purposes. A wide class of visualizations and numerical tests have been included. As future steps we mention
  • fast preconditioning strategies in Krylov methods, when dealing with large quaternion linear systems stemming from modern applications;
  • the definition and study of the GLT $r$-block $d$ level * algebras class in the quaternion setting, $r,d\geq 1$, having in mind variable coefficient problems;
  • quantitative estimates and asymptotic expansions for the spectra of quaternion matrix sequences;
  • extremal spectral and singular value behavior, mimicking the tools and the results in [44, 9, 49, 48, 53];
  • a general analysis beyond the single-axis case;
  • the use of the present powerful machinery in applications and related numerical algorithms.
Regarding the first item, we observe that a notion of linear positive operator can be given in the quaternion setting using the results of the present work. Hence we have the tools for considering the challenge of a Korovkin approximation [30] for the Frobenius optimal preconditioners [46, 47] in fast algebras, following and adapting the technique developed in [46, 47, 12]; see also the Research Project RP 6 in [52]. On a different side, also the topological barriers to the superlinear convergence in the multilevel setting [55, 40] represent an interesting challenge worth investigating.
The second item is very ambitious and requires further study and understanding. Starting from [45], it has been clear that the use of diagonal plus Toeplitz matrix sequences can be very useful for designing fast algorithms in the case of approximated partial differential equation with variable coefficients, when local numerical schemes are used [20, 21, 5]: a big challenge is the quaternion or tensor extensions; see also the Research Project RP 4, items 3, 4, in [52].
With reference to the third item, we remind that linear in time matrix-less algorithms have been developed in [15, 14, 8] for the computation of all the eigenvalues in complex Toeplitz setting, based on proper asymptotic expansions (refer to [8] and references therein). A challenge is to obtain conditions under which a similar type of asymptotic expansions hold in the quaternion case, so allowing the extension of matrix-less algorithms.
Finally, we are convinced that the picture would be complete following the directions indicated in the last three items, having in mind concrete applications and numerical techniques for attacking them.
Acknowledgements
Valerio Loi and Stefano Serra-Capizzano are members of the research group GNCS (Gruppo Nazionale per il Calcolo Scientifico) of INdAM (Istituto Nazionale di Alta Matematica). Furthermore, the third author has been supported by the Theory, Economics and Systems Laboratory (TESLAB) of the Department of Computer Science of the Athens University of Economics and Business and by the PRIN-PNRR project MATHPROCULT (MATHematical tools for predictive maintenance and PROtection of CULTural heritage, Code P20228HZWR, CUP J53D23003780006).